### v1

# Palu: Compressing KV-Cache with Low-Rank Projection

Abstract
KV-Cache compression methods generally sample a KV-Cache of effectual tokens or quantize it into lower bits. However, these methods cannot exploit the redundancy of the hidden dimension of KV tensors. This paper investigates a unique hidden dimension approach called Palu, a novel KV-Cache compression framework that utilizes low-rank projection. Palu decomposes the linear layers into low-rank matrices, caches the smaller intermediate states, and reconstructs the full keys and values on the fly. To improve accuracy, compression rate, and efficiency, Palu further encompasses (1) a medium-grained low-rank decomposition scheme, (2) an efficient rank search algorithm, (3) a low-rank-aware quantization algorithm, and (4) matrix fusion with optimized GPU kernels. Our extensive experiments with popular LLMs show that Palu can compress KV-Cache by more than 91.25% while maintaining a significantly better accuracy (up to 1.19 lower perplexity) than state-of-the-art KV-Cache quantization methods at a similar or even higher memory usage. When compressing KV-Cache for 50%, Palu delivers up to 1.61$\times$ end-to-end speedup for the attention module. Our code is publicly available at: https://github.com/shadowpa0327/Palu.
Palu: Compressing KV-Cache with Low-Rank Projection
[…]
footnotetext: These authors contributed equally to this work.
1 Introduction
Large language models (LLMs) are revolutionizing the AI industry and providing a high-level intelligence that previous machine learning (ML) models could not achieve. To speed inference, caching key-value states (KV-Cache) in memory is a simple yet effective technique. However, the size of the KV-Cache can grow rapidly, straining memory capacity and bandwidth; further, the memory-bounded characteristics of the decoding stage also limit inference speed when loading KV-Cache data. Therefore, KV-Cache compression has become a central research topic for building efficient LLMs.
Figure 1: Palu’s low-rank projection method for KV-Cache reduction. A weight matrix W of linear projection is decomposed into two low-rank matrices. Input X is down-projected to latent representation H, which is cached. Y is reconstructed from H on-the-fly using the up-projection matrix B.
Research on KV-Cache compression falls into two categories: quantization and token eviction. Quantization methods aim to use less bit-width to represent each datum; token eviction techniques seek ways to retain only a partial set of KV-Cache. However, both categories fail to explore the hidden dimensions of KV-Cache where high redundancy often occurs.
This work examines the problem from a fundamentally different perspective, i.e., ways to reduce the hidden dimensions. We introduce Palu, a KV-Cache compression framework that reduces the size of the KV-Cache through low-rank projection. Although low-rank projection has been shown to improve LLM efficiency, e.g., by reducing model size (Yuan et al., 2023; Wang et al., 2024) or minimizing memory footprint for finetuning (Hu et al., 2022; Dettmers et al., 2023), it has not been studied for compressing the KV-Cache, especially in post-training.
A naive way to compress the KV-Cache with low-rank projection is by directly mapping cached matrices into low-rank space (Jolliffe and Cadima, 2016; Zhao et al., 2024). However, this approach imposes an unacceptably heavy overhead of computing the decomposition matrices during runtime, after the prefill stage. To avoid this, Palu statically decomposes the key and value-projection weight matrices and caches the latent representations of the low-rank decomposition (Fig. 1), instead of directly decomposing the KV-Cache. This innovative design enables Palu to reduce memory while substantially mitigating runtime overhead.
Despite its promising memory reduction, we found that naively decomposing the key and value projection matrices along the head dimension causes a substantial loss in accuracy. Although we found that decomposing the entire matrix at one time can maintain accuracy, it significantly increases reconstruction overhead. To find a balance, Palu proposes a medium-grained, group-head low-rank decomposition optimization that is both accurate and less costly in terms of reconstruction.
For LLMs, each linear projection module has a different sensitivity to compression Sharma et al. (2023); Yuan et al. (2023). To exploit the sensitivity and improve accuracy, we design an efficient rank search algorithm based on Fisher information. Our algorithm automatically assigns a larger rank for important matrices and smaller ranks for less critical ones, boosting accuracy at the same compression rate.
In addition to low-rank decomposition, we integrate quantization into Palu. We found that low-rank decomposition often induces severe outliers in the latent representation. Inspired by recent LLM quantization methods Tseng et al. (2024); Ashkboos et al. (2024b), we use Hadamard transformation matrices with a low-rank-aware quantization algorithm to eliminate outliers and increase quantization accuracy. To avoid runtime overhead, Palu fuses the transformation matrices into the forward and backward decomposed matrices, concealing the overhead completely.
To enhance inference latency, we integrate the decomposed matrix into the subsequent linear layer, thereby mitigating the online reconstruction overhead. In scenarios where RoPE is applied and matrix fusion is infeasible, we developed an optimized GPU kernel to accelerate the reconstruction.
We evaluate Palu on popular LLMs and benchmarks. Our experiments show that Palu maintains strong zero-shot accuracy and perplexity with up to 50% low-rank compression. Combining low-rank compression and quantization, Palu compresses KV-Cache by over 91.25% (11.4$\times$ compression) while maintaining perplexity. Compared to a state-of-the-art KV-Cache quantization (Hooper et al., 2024) with 87.5% compression, Palu achieves a significantly lower perplexity of 1.19. Additionally, with 50% compression without quantization, Palu demonstrates a 1.61$\times$ end-to-end speedup, highlighting Palu’s ability to reduce memory usage and enhance speed for LLMs.
Our key contributions include:
  • Palu, a new post-training KV-Cache compression framework that caches low-rank latent representations.
  • Group-head low-rank decomposition (G-LRD), an optimization for balancing accuracy and efficiency.
  • An automated rank search algorithm for adaptively assigning ranks to each decomposed matrix, given a target compression rate.
  • A low-rank-aware quantization algorithm that eliminates low-rank-induced outliers and imposes zero runtime overhead.
2 Background
2.1 Multi-Head Attention Mechanism
The multi-head attention (MHA) mechanism (Vaswani et al., 2017) is a core component of the transformer architecture. Given a new input token $\mathbf{x}\in\mathbb{R}^{d}$, an MHA with $n$ heads projects the input into multiple queries, keys, and values using weight matrices $\mathbf{W}^{q}_{i}$, $\mathbf{W}_{i}^{k}$, and $\mathbf{W}_{i}^{v}$, respectively, for each head $i$, as shown by
$\mathbf{q}_{i}=\mathbf{x}\mathbf{W}_{i}^{q},\hskip 5.0pt\mathbf{k}_{i}=\mathbf{x}\mathbf{W}^{k}_{i},\hskip 5.0pt\mathbf{v}_{i}=\mathbf{x}\mathbf{W}^{v}_{i}.$ (1)
Here, $\mathbf{k}_{i}$ and $\mathbf{v}_{i}$ represent the key and value at time step $t$ for head $i$. We can then compute the attention score for each head $i$ and the corresponding attention output as
$\mathbf{p}_{t,i}=\text{Softmax}\left(\frac{{\mathbf{q}_{i}}\mathbf{K}_{i}^{T}}{\sqrt{d_{h}}}\right),\hskip 5.0pt\mathbf{a}_{i}=\mathbf{p}_{i}\mathbf{V}_{i},$ (2)
where $\mathbf{K}_{i}$ and $\mathbf{V}_{i}$ denote the concatenation of current and all previous keys and values corresponding to the $i$-th head. The final MHA output is obtained by concatenating the outputs of all heads and then applying the out-projection layer $\mathbf{W}_{o}$, as shown by
$\text{MHA}(\mathbf{x})=\sum_{i=1}^{h}\mathbf{a}_{i}\mathbf{W}^{o}_{i}=\sum_{i=1}^{h}(\mathbf{p}_{i}\mathbf{V}_{i})\mathbf{W}^{o}_{i},$ (3)
where $\mathbf{W}^{o}_{i}\in\mathbb{R}^{d_{h}\times d}$ represents the submatrices of the out-projection matrix for each head $i$.
2.2 Singular Value Decomposition (SVD)
SVD (Jolliffe and Cadima, 2016) is a commonly used technique for computing the low-rank approximation for a given matrix. We now introduce our specific use of SVD for Palu’s default low-rank decomposition method.
Given a weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, SVD decomposes $\mathbf{W}$ into three matrices: $\mathbf{W}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{T}$. Here, $\mathbf{U}$ and $\mathbf{V}$ are orthogonal matrices containing the left and right singular vectors, respectively. The matrix $\boldsymbol{\Sigma}$ is a diagonal matrix that consists of singular values. We describe the decomposition as
$\mathbf{W}\approx\mathbf{A}\mathbf{B},\hskip 5.0pt\mathbf{A}=\mathbf{U}_{r}\sqrt{\boldsymbol{\Sigma}_{r}},\hskip 5.0pt\mathbf{B}=\sqrt{\boldsymbol{\Sigma}_{r}}\mathbf{V}_{r}^{T},$
where $\mathbf{A}\in\mathbb{R}^{m\times r}$, $\mathbf{B}\in\mathbb{R}^{r\times n}$, $\boldsymbol{\Sigma}_{r}\in\mathbb{R}^{r\times r}$. $\boldsymbol{\Sigma}_{r}$ is a diagonal matrix containing the largest $r$ singular values, and $\mathbf{U}_{r}$, $\mathbf{V}_{r}^{T}$ are corresponding singular vectors truncated from $\mathbf{U}$ and $\mathbf{V}^{T}$. This truncation and subsequent matrix formation let us approximate weight matrix $\mathbf{W}$ with two low-rank matrices $\mathbf{A}$ and $\mathbf{B}$, thereby reducing the storage by $\frac{mr+rn}{mn}$.
Figure 2: Palu uses low-rank decomposition $(\mathbf{W}\approx\mathbf{A}\mathbf{B})$ to project the key (or value) to a lower-dimensional latent representation $(\mathbf{h})$, thereby reducing the size of the KV-Cache. The original key $(\mathbf{K}_{t})$ is reconstructed on-the-fly with $\mathbf{B}^{k}$, and $\mathbf{B}^{v}$ is fused into $\mathbf{W}^{o}$ to avoid reconstruction overhead. The fusion also reduces the computational burden for output projection.
3 The Palu Framework
3.1 Compressing the KV-Cache via Low-Rank Projection
As noted, a naive approach to compress the KV-Cache with low-rank projection is to apply SVD directly on the KV-Cache and store the top-$r$ singular vectors. However, this approach poses significant computational challenges during runtime that make it impractical for deployments (see Section 1).
To apply low-rank projection more efficiently, Palu uses SVD to decompose the key and value projection weight matrices. This approach is based on the observation that low-rank decomposition rewrites the linear projection layer from $\mathbf{y}=\mathbf{x}\mathbf{W}$ into $\mathbf{y}=\mathbf{x}\mathbf{A}\mathbf{B}$. Here, $\mathbf{A}\in\mathbb{R}^{d\times r}$ is the low-rank projection matrix, and $\mathbf{B}\in\mathbb{R}^{r\times d}$ is the reconstruction matrix derived by SVD. The forward process first down-projects the input token $\mathbf{x}\in\mathbb{R}^{d}$ into a low-dimensional latent space $\mathbf{h}\in\mathbb{R}^{r}$ and then up-projects it back to the original space:
$\mathbf{h}=\mathbf{A}\mathbf{x},\quad\mathbf{y}=\mathbf{B}\mathbf{h}$ (4)
This two-step process lets Palu (1) store the lower dimension latent representation instead of the origin keys and values states, and (2) reconstruct them during decoding.
3.1.1 Integration with the Attention Mechanism and Offline Matrix Fusion
We now describe how Palu decomposes the key and value linear layers for the attention mechanism. For each attention head $i$, Palu applies SVD and maps the key-projection matrix $\mathbf{W}^{k}_{i}$ and value-projection matrix $\mathbf{W}^{v}_{i}$ into $\mathbf{A}^{k}_{i}\mathbf{B}^{k}_{i}$ and $\mathbf{A}^{v}_{i}\mathbf{B}^{v}_{i}$.
Based on the formula of attention output in equation 2, Palu absorbs the reconstruction matrix $\mathbf{B}^{v}_{i}$ into the output projection matrix $\mathbf{W}_{i}^{o}$ offline:
$\mathbf{a}_{i}\mathbf{W}^{o}_{i}$ $=(\mathbf{p}_{i}\mathbf{V}_{i})\mathbf{W}^{o}_{i}$ (5)
$=(\mathbf{p}_{i}\mathbf{H}^{v}_{i}\mathbf{B}^{v}_{i})\mathbf{W}^{o}_{i}$
$=\mathbf{p}_{i}\mathbf{H}^{v}_{i}(\mathbf{B}^{v}_{i}\mathbf{W}^{o}_{i})$
Such fusion lets Palu skip the explicit reconstruction of the full value vectors, reduce the number of matrix multiplications, and improve efficiency. A similar approach applies for calculating attention scores. Matrix $\mathbf{B}^{k}_{i}$ can be fused into the query projection matrix $\mathbf{W}_{i}^{q}$ offline, as shown by
$\mathbf{q}_{i}\mathbf{K}_{i}^{T}$ $=\mathbf{q}_{i}(\mathbf{H}^{k}_{i}\mathbf{B}^{k}_{i})^{T}$ (6)
$=\mathbf{x}_{t}\mathbf{W}^{q}_{i}(\mathbf{B}^{k}_{i})^{T}(\mathbf{H}^{k}_{i})^{T}$
$=\mathbf{x}_{t}\Bigl{(}\mathbf{W}^{q}_{i}(\mathbf{B}^{k}_{i})^{T}\Bigr{)}(\mathbf{H}^{k}_{i})^{T}.$
Here, $\mathbf{B}^{k}_{i}\in\mathbb{R}^{r\times d_{h}}$ and $\mathbf{W}^{q}_{i}\in\mathbb{R}^{d\times d_{h}}$, so the fused matrix $(\mathbf{W}^{q}_{i}(\mathbf{B}^{k}_{i})^{T})$ has size $\mathbb{R}^{d\times r}$. This fusion boosts computational efficiency by reducing the matrix dimension during attention score calculation.
3.1.2 Mechanism with Positional Embedding
Recent LLMs, such as the Llama family, apply positional embedding (i.e., RoPE (Su et al., 2021)) onto the query and key states before multiplication. The non-linear operations of adding positional embedding prevent the fusion for the attention score, as described in equation 6. In this case, Palu reconstructs the keys from latent representations on-the-fly during decoding. We improve reconstruction efficiency by designing an optimized kernel, which we present in Sec. 3.5.
Note that for some positional embedding methods, such as ALiBi Press et al. (2022), or new attention mechanisms, such as MLA DeepSeek-AI et al. (2024), positional embedding is not directly applied on the key states; therefore, the fusion of equation 6 remains valid. We expect the end-to-end speedup to be larger on these models. Fig. 2 presents a detailed workflow of Palu based on Llama’s attention with RoPE.
3.2 Decomposition Granularity
3.2.1 Multi-Head Low-Rank Decomposition
We name the per-head decomposition scheme in Sec.3.1.1 as multi-head low-rank decomposition (M-LRD). We found M-LRD often causes a non-negligible accuracy degradation (discussed further in Sec.4.1), possibly because SVD fails to capture the common information shared across heads. Therefore, alternative approaches are needed for preserving model accuracy.
3.2.2 Joint-Head Low-Rank Decomposition
An alternative approach is to jointly decompose weight matrices for all heads. By considering the combined weight matrix $\mathbf{W}_{\text{joint}}=[\mathbf{W}_{1},\mathbf{W}_{2},\dots,\mathbf{W}_{n}]\in\mathbb{R}^{d\times(d_{h}\cdot n_{h})}$, we can perform a single low-rank decomposition $\mathbf{W}_{\text{joint}}\approx\mathbf{A}_{\text{joint}}\mathbf{B}_{\text{joint}}$, where $\mathbf{A}_{\text{joint}}\in\mathbb{R}^{d\times r_{\text{joint}}}$ and $\mathbf{B}_{\text{joint}}\in\mathbb{R}^{r_{\text{joint}}\times(d_{h}\cdot n_{h})}$. We call this scheme joint-head low-rank decomposition (J-LRD).
J-LRD has the advantage of preserving the common principal components shared among different heads. This occurs because SVD is particularly effective at capturing the dominant components when applied to a larger, combined matrix, resulting in a more accurate approximation.
For J-LRD, the joint latent representation shared among all heads can be computed with $\mathbf{h}_{\text{joint}}=\mathbf{x}\mathbf{A}_{\text{joint}}$. During decoding, the original states for each head can be reconstructed via
$\bigl{[}\mathbf{y}_{1},\dots,\mathbf{y}_{n}\bigr{]}$ $=\mathbf{h}_{\text{joint}}\mathbf{B}_{\text{joint}}.$
High Inference Overhead.
Despite better preserving model information, J-LRD introduces significant computational and memory overhead during decoding. Specifically, the total number of floating point operations (FLOPs) to reconstruct the key or values state of one head now becomes $r_{\text{joint}}\cdot d_{h}\cdot n$. Assuming the same size as the total low-rank latent representations (i.e., $r_{\text{joint}}=\sum_{i=1}^{n}{r_{i}}$), the total reconstruction cost is $n$ times higher than M-LRD, whose total FLOPs is $r_{i}\cdot d_{h}\cdot n$. When considering the matrix fusion in Sec.3.1.1, the fused matrix size of J-LRD has a size of $r_{joint}\cdot d\cdot n$, which is also $n$ times larger than M-LRD, leading to a substantial higher memory consumption.
Figure 3: Performing decomposition at different granularities. Jointly decomposing multiple heads can achieve higher accuracy. Assuming the same total size of the latent representations (i.e., $4\cdot r_{i}=2\cdot r_{g}=r_{\text{joint}}$), the FLOPs for reconstruction overhead in joint-head decomposition schemes are 4 times larger than those in multi-head ones.
3.2.3 Group-Head Low-Rank Decomposition
To balance the trade-off between accuracy and reconstruction cost, we propose group-head low-rank decomposition(G-LRD). G-LRD decomposes the matrices for a group of heads together. With combined weight matrices, it captures shared information within each group while limiting computational overhead and preserving accuracy.
To illustrate the G-LRD process, consider the weight matrices for a group of $s$ heads, $\mathbf{W}_{g_{j}}=\bigl{[}\mathbf{W}_{j,1}\dots\mathbf{W}_{j,s}\bigr{]}$, where $\mathbf{W}_{g_{j}}\in\mathbb{R}^{d\times(d_{h}\cdot s)}$. We low-rank decompose $\mathbf{W}_{g_{j}}\approx\mathbf{A}_{g_{j}}\mathbf{B}_{g_{j}}$, where $\mathbf{A}_{g_{j}}\in\mathbb{R}^{d\times r_{g}}$ and $\mathbf{B}_{g_{j}}\in\mathbb{R}^{r_{g}\times(d_{h}\cdot s)}$. The latent representation shared among attention heads in the same group can be computed as $\mathbf{h}_{\text{g}_{j}}=\mathbf{x}\mathbf{U}_{\text{g}_{j}}$. During decoding, the original key or value for each head can be reconstructed via
$[\mathbf{y}_{j,1}\dots\mathbf{y}_{j,s}]=\mathbf{h}_{\text{g}_{j}}\mathbf{B}_{\text{g}_{j}}.$
The FLOPs for reconstructing the keys and values for each head in G-LRD is $r_{g}\cdot d_{h}\cdot n_{g}$, where $n_{g}=\frac{n}{s}$ is the number of groups. Comparing the cost to J-LRD and assuming the same total rank size ($r_{g}\cdot n_{g}=r_{\text{joint}}$), G-LRD reduces the reconstruction cost by $n_{g}$. Similarly, G-LRD also reduces the fused matrix size by $n_{g}$. To sum up, G-LRD offers a middle ground between computation overhead and approximation accuracy. We illustrate M-LRD, J-LRD and G-LRD in Fig.3.
For further discussion on decomposed weight matrix size, please refer to Appendix B, where more detailed information is provided.
3.3 Automatic Rank Allocation
To allocate an ideal rank size to the decomposition target, it is crucial to accurately estimate the importance of the target matrix (e.g., grouped weights). In Palu, we identify Fisher information (Ly et al., 2017; Liu et al., 2021) as an accurate approximator since it can quantify the amount of information for each parameter. We then employ the sum of Fisher information to estimate the importance of the weight matrix of each linear layer.
Assuming that the compression sensitivity is proportional to Fisher information, we determine the rank for each weight matrix by computing the ratio of its Fisher information to the total Fisher information across all decomposition targets. We use this ratio to allocate the compression rate (i.e., rank level $r$), ensuring that more important layers retain higher rank levels.
For a detailed ablation study on our Automatic Rank Allocation method, please refer to Appendix H.3, where we analyze the impact of our strategy on model performance.
3.4 Low-Rank-Aware Quantization
Figure 4: Activation distribution of the low-rank compressed key caches at the $4^{th}$ Llama-2 attention layer.
We integrate quantization into Palu to further compress the KV-Cache. However, we observe that low-rank compressed latent representations have severe outliers, which limit quantization applicability in Palu. Unlike natural outliers described in previous KV-Cache quantization literature (Liu et al., 2024; Hooper et al., 2024), these outliers are induced by SVD-based low-rank factorization.
Fig. 4 (a) shows the distribution of low-rank compressed key states from a layer of Llama-2 with G-LRD. Repeating outlier patterns appear at the beginning of each decomposed group because SVD arranges larger eigenvalues in the initial rows or columns, resulting in rapidly descending values in the latent representation. This pattern stretches the data distribution and hurts quantization accuracy.
Inspired by recent LLM quantization literature (Ashkboos et al., 2024b; Tseng et al., 2024), we apply the Walsh-Hadamard transform (WHT, Fino and Algazi) to eliminate outliers (Fig. 4 (b)), enabling a high quantization accuracy. However, this transformation introduces an extra matrix multiplication with associated runtime overhead. Unlike earlier methods (Ashkboos et al., 2024b) that must apply online WHT when quantizing KV-Cache, we optimize this process by integrating the Hadamard matrix into low-rank decomposed weights with no additional overhead, as described by
$\mathbf{W}\approx\mathbf{AB}=(\mathbf{AR})(\mathbf{R}^{T}\mathbf{B})=\mathbf{\hat{A}\hat{B}},$ (7)
where $\mathbf{R}$ is the Hadamard matrix. This optimization allows Palu can integrate the proposed low-rank compression technique with low-bit quantization. Our experiments show that, on top of the low-rank compression, our quantization method only negligibly increases perplexity, even at extreme levels such as 3-bit or 2-bit (see Sec.4.2).
3.5 Efficient Inference Implementation
Support for RoPE-based Attention.
To support RoPE-based attention, we fuse the reconstruction of original keys, applying positional embedding (i.e., RoPE (Su et al., 2024)) and the final matrix-vector multiplication of query and reconstructed keys by implementing customized GPU kernels in Triton Tillet et al. (2019). We leverage the tiling techniques akin to FlashAttention (Dao et al., 2022) to minimize the memory reads/writes of GPU high bandwidth memory.
Table 1: Perplexity and zero-shot accuracy of Palu, with different decomposition strategies at 50% compression rate.
Model Method Perplexity $\downarrow$ Zero-Shot Accuracy (%) $\uparrow$
Wiki2 C4 OBQA Hella PIQA ARC-e ARC-c Wino Avg.
Llama2-7B Baseline 5.47 7.26 44.20 76.00 78.07 76.30 46.42 69.30 65.05
J-LRD 5.62 7.75 45.40 75.57 77.48 75.97 45.31 69.22 64.82
G-LRD 6.01 9.82 43.60 73.39 76.33 73.02 42.57 66.77 62.61
M-LRD 6.75 12.01 39.60 65.35 74.76 67.17 35.24 64.64 57.79
Llama2-13B Baseline 4.88 6.70 45.20 79.39 79.11 79.42 49.06 72.38 67.43
J-LRD 4.97 6.92 46.40 79.48 78.62 79.29 49.91 70.56 67.38
G-LRD 5.31 7.76 45.60 77.29 77.42 76.05 45.99 72.45 65.80
M-LRD 5.65 8.34 43.20 74.34 77.53 75.76 45.39 68.98 64.20
4 Experiments
Table 2: Accuracy of Palu with 30% and 50% compression rate on 8 tasks in LongBench.
Model Method SAMSum LCC M-News Qasper TREC QmSum R-Bench TriQA Avg.
Vicuna-7B Baseline 40.73 49.20 27.98 25.57 69.00 22.54 43.50 86.42 45.62
Palu-30% 40.42 49.00 26.96 18.54 65.50 21.78 45.77 85.22 44.15
Palu-50% 38.36 34.59 24.78 19.77 63.50 21.44 35.42 79.74 39.70
Mistral-7B Baseline 41.71 66.59 20.26 8.13 67.00 19.99 59.10 89.90 46.59
Palu-30% 43.23 64.23 23.50 10.21 72.50 21.16 56.80 89.82 47.68
Palu-50% 36.43 50.22 19.23 7.54 67.50 19.56 44.76 85.37 41.33
Models and Tasks.
We evaluate Palu on three LLM families, Llama2(Touvron et al., 2023), Mistral(Jiang et al., 2023) and Vicuna(Chiang et al., 2023). For accuracy evaluation, we measure perplexity on the WikiText-2 (Merity et al., 2016) and C4 (Raffel et al., 2020) datasets and use LM-Evaluation-Harness (Gao et al., 2023) to measure zero-shot accuracy on six common sense tasks. We also evaluate long context accuracy on 8 tasks in LongBench (Bai et al., 2023). See Appendix I for further details on the dataset and settings.
Compression Settings.
We implemented Palu using the Huggingface library Wolf et al. (2020). Decomposition of the key and value projection layers was performed using the truncation-aware SVD method proposed by SVD-LLM (Wang et al., 2024). Unless otherwise specified, results for Palu are reported using G-LRD with a group size (gs) of 4 with equal rank size for each group. To calculate Fisher information in rank searching, we used 2048 samples from Wikitext-2, each with a sequence length of 1024. For quantization, we aligned our settings with Atom (Zhao et al., 2023), and use a simple per-token, asymmetric quantization to quantize the low-rank latent representation of both key and value states.
4.1 Main Results
Perplexity Evaluation.
We first present perplexity evaluation results on Wikitext-2 and C4 in Table 1. For the Llama2-7B model, we observe that the M-LRD method fails to maintain low perplexity at a 50% compression rate. In contrast, the J-LRD method performs significantly better given the same compression rate, achieving perplexity values of 5.62 on Wikitext-2. The G-LRD method, with a group size of 4, achieves an optimal balance between perplexity and computational cost, yielding perplexity values of 6.01 on Wikitext-2 at a 50% compression rate. The same trend is consistent across the Llama2-13B model.
Zero-shot Evaluation Results.
We further evaluate Palu on six common-sense zero-shot tasks, as shown in Table 1. Similar to the perplexity evaluation, the J-LRD method demonstrates the best performance on Llama2-7B, with only a 0.23% average accuracy degradation, while the M-LRD method results in the lowest average performance, with a 7.26% drop in accuracy compared to the baseline. In comparison, G-LRD experiences only a 2.4% average accuracy decline.
LongBench Results.
To assess whether Palu can effectively compress the KV-Cache for long-context scenarios, we present the performance of Palu across eight tasks from LongBench. We select Mistral-7B and Vicuna-7B (Chiang et al., 2023) as our evaluation targets, which support sequences up to 8k and 16k, respectively. Table 2 reports experiment results. We find that at a 50% compression level, it is relatively difficult to fully preserve accuracy. However, when compressing KV-Cache for 30%, Palu only has a 1.47% averaged accuracy degradation for Vicuna-7B, and even achieves a 1.09% enhancement for Mistral-7B. These results evidence of Palu’s ability to shrink the KV-Cache while maintaining performance.
For detailed information on the integration of Palu with quantization and the performance measurements on LongBench, please refer to Appendix E.
4.2 Integrating Palu with Quantization
Table 3 shows the results of integrating quantization with Palu. With 3-bit quantization, Palu exhibits only a negligible perplexity increase of 0.08 at a 30% compression rate and a minimal 0.23 at a 50% compression rate, both compared to Palu at 16-bit. This indicates a minimal trade-off for substantial compression of integrating quantization. Furthermore, at 2-bit quantization, Palu significantly outperforms KVQuant, which is a SOTA quantization-based method. Palu demonstrates a compelling reduction in perplexity by 1.19 and 0.54 points, while consuming 30% and 50% less memory, respectively. These results underscore the effectiveness of Palu, establishing it as a superior KV-Cache compression strategy.
4.3 Latency Evaluation
We now provide the latency evaluation for Palu. In our experiment, latency is measured on an A6000 GPU, and the averaged results over 100 runs are reported. For baseline, we use the default implementation from HuggingFace in PyTorch. We do not apply FlashAttention (Dao et al., 2022) for an apple-to-apple comparison, as our current implementation does not fuse the Softmax operation. We leave the integration with FlashAttention as the future work.
GEMV with Online Reconstruction.
We evaluate the performance of our fused reconstruction kernel (Sec. 3.5) for computing RoPE-based attention score. Specifically, we measure the latency from pre-RoPE query vector to the post-GEMV attention score (refer to Fig. 2 for details). Here, we set the compression rate for our kernel as 75%. It’s because we observe the key states often achieve such compression rate when setting a 50% overall compression rate (see Appendix H.3 for detailed distribution).
We compare Palu’s runtime against the baseline, which is simply a GEMV, and report the speedup in Fig. 5 (a). For group size 32 (J-LRD), although it’s the most accurate decomposition, the high reconstruction leads to a significant 7.09$\times$ slowdown at sequence length 64K . For group size 1 (M-LRD), our fused kernel achieves up to a 2.48$\times$ speedup at a sequence length 16k, showcasing a high speedup when more accuracy loss is acceptable. For group size 2 and 4 (G-LRD), our fused kernel demonstrates 1.75$\times$ and 1.16$\times$ speedup at sequence length 16K. These results highlight the importance of exploring trade-off between decomposition granularity, reconstruction cost, and accuracy. Furthermore, we observe our kernel speedup decreases for sequence lengths over 16k, due to the increasing reconstruction cost shifts the attention operation from memory- to compute-bounded.
Table 3: Perplexity and KV-Cache size comparison of Palu and other quantization-based methods for Llama2-7B on WikiText-2. For perplexity, sequence length is set to 4096, aligning with baselines. KV-Cache size is demonstrated for a sequence length of 128K.
Method Bit PPL
KV-Cache
Size (GB)
Comp.
Rate
Baseline 16 5.12 64.0 -
Palu-30% 16 5.25 44.8 30%
Palu-50% 16 5.63 32.0 50%
Atom 3 6.15 12.6 80.32%
KVQuant 3 5.35 12.0 81.25%
Palu-30% 3 5.33 8.4 86.87%
Palu-50% 3 5.77 6.0 90.63%
Atom 2 117.88 8.6 86.56%
KVQuant 2 6.95 8.0 87.50%
Palu-30% 2 5.76 5.6 91.25%
Palu-50% 2 6.41 4.0 93.75%
End-to-end attention module.
We measure the end-to-end latency of the entire RoPE-based attention module, with the integration of our fused reconstruction kernel for key states (Sec. 3.5) and offline matrix fusion for value states (Eq. 5). Here, we set the compression rate for key and value as 75% and 25%, for an overall 50% compression rate. The detailed forward pass operations are presented in Fig. 2. We measure the decoding runtime for sequence length from 4K to 64K, and report the runtime of Palu and baseline in Fig. 5 (b). As Fig. 5 (b), Palu with smaller group sizes consistently delivers more than 1.43$\times$ speedup across all sequence lengths. For group size 1 (M-LRD), Palu achieves a 1.98$\times$, close to optimal (for 50% compression rate), speedup for sequence length 16K. For group size 4 (G-LRD), which is the sweet spot for maintaining strong accuracy and limiting reconstruction cost, Palu achieves up to 1.61$\times$ end-to-end speedup at sequence length 64K.
We observe increasing speedup as sequence length increases from 4k to 64k due to continuously reducing latency from compressing value states. However, speedup is expected to converge or slightly decrease with further length increases due to higher reconstruction costs and the compute-bound nature of reconstruction kernels. We expect the speedup of Palu will be larger for non-RoPE-based attention, as the reconstruction overhead can be entirely eliminated by applying Eq. 6.
For the end-to-end model latency, please refer to Appendix D, where more detailed information and data are provided.
Figure 5: Normalized speedup for (a) reconstruction kernel at 75% compression and (b) RoPE-based attention module with an overall 50% compression under different group sizes.
5 Related Work
SVD for LLM Compression.
There are several related works applying SVD for LLM compression. An early work (Noach and Goldberg, 2020) simply compresses the weight matrices using standard SVD, resulting in high compression errors. To address this, FWSVD (Hsu et al., 2022) incorporates Fisher information to prioritize parameters. Additionally, ASVD (Yuan et al., 2023) taken the activation outlier phenomenon in LLM into consideration, and propose and scaling techniques to mitigate the compression error. SVD-LLM (Wang et al., 2024) further advances this work by understanding the compression loss from each singular value, improving compression accuracy. Unlike previous related SVD works that focus on compressing model weights, Palu focus on using SVD to reduce KV-Cache size.
KV-Cache Quantization.
Quantization appears to be an effective compression technique to compress KV-Cache. Atom (Zhao et al., 2023) uses a simple per-token quantization scheme to compress KV-Cache. WKVQuant (Yue et al., 2024) proposes a two-level quantization scheme to improve quantization accuracy. KIVI (Liu et al., 2024) utilizes different quantization schemes, per-channel and per-token, for key and value states respectively. KVQuant (Hooper et al., 2024) follow a similar setting but adopts non-uniform quantization and sparse matrix for preserving outliers. In Palu, we eliminate outliers with WHT to enhance quantization friendliness and, therefore, only need a simple per-token quantization and achieve a remarkable accuracy.
6 Conclusion
We introduce Palu, a novel KV-Cache compression framework that decomposes key and value linear projection weight matrices and caches the compressed latent representations. We propose various optimizations, including group-head low-rank decomposition, rank search algorithm, low-rank-aware quantization, matrix fusion, and kernel fusion. With these optimizations, Palu can maintain accuracy while achieving significant memory reduction and faster inference. Experiments show that Palu can accelerate the end-to-end attention module by up to 1.61$\times$ with 50% KV-Cache compression; when combined with 2-bit quantization, Palu achieves a remarkable 91.25% compression rate, both maintaining strong accuracy.
Limitations and Future Work
We introduce a novel KV-Cache compression framework based on low-rank projection. However, our experiments cover model variants with up to only 13B parameters due to limited computational resources. Future work needs to establish how Palu performs for models with larger sizes (e.g., 70B or beyond). Also, we leave for future work the evaluation of end-to-end speedup when combining low-rank compression and quantization. It would also be interesting to investigate how to combine Palu with other efficient LLM methods, such as token eviction or weight quantization, to achieve the highest possible acceleration.
Acknowledgement
We would like to thank Sandy Kaplan for her valuable feedback and support in refining the writing of this paper. Her insights and suggestions significantly contributed to the clarity and quality of the final manuscript.
Appendix
Appendix A Quantization Basics
Quantization techniques use discrete low-bit values to approximate high-precision floating points. The general asymmetric uniform quantization function is defined as:
$\overline{\mathbf{X}}=\text{clamp}\Big{(}\left\lfloor\frac{\mathbf{X}}{s}\right\rceil+z,0,2^{B}-1),$ (8)
[…]
Appendix B Discussion regarding memory usage
In this work, the experimental results focus on the compression rate of the KV-Cache as a key metric. However, it is crucial to consider overall memory savings as a more significant factor. For instance, as demonstrated in Sec. 2.2, a typical compression rate of 30% can lead to an increase in weight size by approximately 40%. This increase is calculated under the assumption that $m=n$ and $r=0.7n$, resulting in the equation $\frac{mr+nr}{mn}=1.4$. Such an increase indicates substantial extra memory usage.
This issue primarily arises in J-LRD decomposition schemes, where the projections of all heads are decomposed jointly. In contrast, our M-LRD decomposition schemes and optimized G-LRD schemes involve non-square target matrices. For example, in the G-LRD scheme with a group size of 4, the target matrix is formed by concatenating the original projection matrices of each attention head in the group. In the Llama-2-7b model, with an embedding dimension of 4096 and head dimensions of 128, each projection matrix is 4096x128, resulting in a concatenated matrix of size 4096x512. In this case, the dimensions should be considered as $m=8n$. Applying the referenced equation $\frac{mr+nr}{mn}$ with $r=0.7n$, we find that $\frac{mr+nr}{mn}=0.7875$, indicating no additional storage cost and, in fact, achieving an additional 21.25% memory savings.
Furthermore, it is important to highlight that the weights associated with the K and V projections account for only 2 out of 7 linear layers within transformer blocks, comprising merely 16% of the parameters in Llama-2-7b models. This limits the overall impact on memory usage. Thus, while J-LRD may incur overhead, the M-LRD and G-LRD schemes provide efficient alternatives that do not lead to increased memory usage, making them viable options for practical applications.
Appendix C Overall Memory Reduction
In Fig. 6, we show the total memory usage, including model weights and KV-Cache under different sequence lengths. We observe that KV-Cache quickly becomes the memory bottleneck as sequence length grows. At a sequence length of 64k, the KV-Cache constitutes 78% of the total memory consumption. With 50% low-rank compression, Palu effectively reduces memory usage by 1.7$\times$, and when combined with 2-bit quantization, it further decreases total memory usage by 4.6$\times$.
Figure 6: Total memory usage for various sequence lengths in Llama2-7B. low-rank compression is at 50% for Palu.
Appendix D End-to-End Latency Measurements
We provide the latency evaluation of integrating our optimized attention modules into the Llama-2 models in Table 4. Measurement times are expressed in milliseconds (ms). Our measurements indicated that Palu achieved significant end-to-end speedup during the decoding phase, especially with longer sequence lengths. Specifically, for a sequence length of 32K, the speedups were 1.40x, 1.55x, and 1.71x for gs=4, gs=2, and gs=1, respectively. The results demonstrate the effectiveness of our methods and the potential to explore the redundancy in the hidden dimension.
Table 4: End-to-end decoding latency of Llama-2-7B with Palu measured on A6000.
Method Group Sequence Length
Size 4k 8k 16k 32k
Baseline - 36.07 44.03 69.28 112.75
J-LRD 32 86.39 103.41 140.85 OOM
G-LRD 4 32.12 38.89 52.69 80.81
G-LRD 2 28.82 34.94 47.45 72.89
M-LRD 1 27.00 32.46 43.56 66.13
Appendix E Extended Experiments on LongBench and comparisons to quantization baselines
We conducted extensive experiments using the Mistral-Instruct-7B-v0.2 (Jiang et al., 2023) and LongChat-7b-v1.5-32k (Li et al., 2023) models, both with extended context lengths of 32k tokens. We included 16 available English datasets from LongBench to ensure a comprehensive evaluation. We report the average score for each type of task separately, as well as the overall average across all 16 tasks. This evaluation aimed to broadly assess the effectiveness of Palu in long-context scenarios. For reference, we also include comparisons with KIVI.
Our findings indicate that at a 30% compression level, Palu achieves only a minor average accuracy drop (<1%) compared to the baseline for both models. These impressive results are attributed to our effective use of low-rank projection methods to explore redundancy in hidden dimensions. Furthermore, by fusing the Hadamard matrix offline to improve quantization friendliness, we were able to quantize the low-rank projected KV-Cache down to 3 bits. This quantization process resulted in less than a 1% accuracy degradation, achieved with a straightforward per-token quantization approach. When compared to KIVI, our framework demonstrates the ability to harness an additional 30% of redundancy. Notably, Palu does not require the complex grouped quantization and mixed-precision techniques employed by KIVI, highlighting the quantization friendliness that Palu offers.
Table 5: Experiment results on Longbench. Average bit widths are estimated for each approach, assuming 10K context length. Each column represents the average score for the tasks of each type. Multi-QA and Single-QA denote Multi-Document and Single-Document Question Answering, respectively, while Summ refers to Summarization.
Model Method Avg. Comp. Multi- Single- Summ Few-Shot Code Synthetic Avg.
Bits Ratio QA QA
Mistral-7B-v0.2 Baseline 16 1.00x 29.63 36.43 28.10 66.71 54.16 44.87 42.54
Palu-30% 16 1.43x 29.83 36.52 27.48 65.70 55.16 37.92 41.55
Palu-50% 16 2.00x 26.92 35.33 26.01 64.04 44.54 16.88 36.23
KIVI-2 3.16 5.05x 28.81 35.07 27.60 66.45 54.47 40.28 41.45
Palu-30% (3 bits) 3 7.59x 29.48 36.40 27.20 65.73 53.19 34.74 40.77
Palu-50% (3 bits) 3 10.6x 26.73 32.72 25.73 63.25 44.43 18.57 35.71
LongChat-7B-v1.5 Baseline 16 1.00x 23.95 31.12 26.74 63.80 56.91 15.25 36.32
Palu-30% 16 1.43x 22.42 29.43 25.52 62.87 58.99 14.25 35.45
Palu-50% 16 2.00x 22.61 25.33 22.73 60.12 43.52 6.84 30.82
KIVI-2 3.16 5.06x 23.24 30.19 26.47 63.54 53.51 16.13 35.60
Palu-30% (3 bits) 3 7.59x 23.12 29.21 25.04 61.99 54.38 11.25 34.33
Palu-50% (3 bits) 3 10.6x 18.56 24.14 22.35 58.76 40.50 6.02 29.03
Table 6: Evaluation integrating LoRA with Palu on Llama2-7B.
Comp. Rate Method Zero-Shot Accuracy (%) $\uparrow$
OBQA HellaSwag PIQA ARC-e ARC-c WinoGrande Avg. Diff.
Rate = 0% baseline 44.20 76.00 78.07 76.30 46.42 69.30 65.05 -
Rate = 50% w/o LoRA J-LRD 45.40 75.57 77.48 75.97 45.31 69.22 64.83 -0.22
G-LRD 43.60 73.39 76.33 73.02 42.57 66.77 62.61 -2.44
M-LRD 39.60 65.35 74.76 67.17 35.24 64.64 57.79 -7.26
Rate = 50% w/ LoRA J-LRD 44.20 74.09 78.51 77.27 48.81 71.03 65.65 +0.60
G-LRD 43.40 73.08 78.56 75.72 47.10 69.85 64.62 -0.43
M-LRD 41.80 70.78 78.02 73.86 43.86 69.22 62.92 -2.12
Table 7: Zero-shot accuracy of Palu on LLama2 models, with different decomposition strategies at 30% compression rate.
Model Method Zero-Shot Accuracy (%) $\uparrow$
OBQA Hella PIQA ARC-e ARC-c Wino Avg. Diff.
Llama2-7B Baseline 44.20 76.00 78.07 76.30 46.42 69.30 65.05 -
J-LRD 44.40 76.01 77.75 76.01 45.31 69.38 64.81 -0.24
G-LRD 45.20 76.07 78.56 75.93 44.97 68.75 64.91 -0.14
M-LRD 43.20 73.50 77.42 74.16 42.58 68.90 63.29 -1.76
Llama2-13B Baseline 45.20 79.39 79.11 79.42 49.06 72.38 67.43 -
J-LRD 45.20 79.58 80.57 79.67 49.40 71.74 67.69 +0.26
G-LRD 46.60 79.04 78.23 79.42 48.38 71.59 67.21 -0.22
M-LRD 45.40 78.57 78.78 78.83 47.61 72.06 66.88 -0.55
Appendix F Integrating Palu with LoRA Finetune
LoRA (Hu et al., 2022) has become one of the most widely used efficient fine-tuning techniques for adapting models to particular tasks or domains with limited data. It has also been applied with LLM compression approaches (Wang et al., 2024; Ma et al., 2023) as a post-compression recovery technique to recover information loss after compression. In Palu, LoRA is also applicable to boost the accuracy further.
To integrate LoRA with Palu, we introduce additional low-rank matrices $\mathbf{A}_{r}^{\prime}\in\mathbb{R}^{d\times r^{\prime}}$ and $\mathbf{B}_{r}^{\prime}\in\mathbb{R}^{r^{\prime}\times d}$ to refine the original low-rank projection as below:
$\mathbf{h}=\mathbf{A}\mathbf{x}+\mathbf{A}_{\mathbf{r^{\prime}}}\mathbf{B}_{\mathbf{r^{\prime}}}\mathbf{x}$ (9)
Here, $\mathbf{A}$ will be fixed parameters derived from low-rank decomposition from pre-trained weights of linear layers, while $\mathbf{A}_{r}^{\prime}$ and $\mathbf{B}_{r}^{\prime}$ are trainable parameters to capture the task-specific nuances and recovers the information lost during the compression.
Setup.
Following Ashkboos et al. 2024a, we sample 8k samples from the Alpaca training dataset as a fine-tuning dataset and apply LoRA with rank $r^{\prime}=32$ and $\alpha=32$. All other hyper-parameters are aligned with Ashkboos et al. (2024a), except for the learning rate $42e-$ 4, and the use of a cosine learning rate scheduler.
Experiment Results.
We present the experiment results with LoRA in Table 6. Following Ashkboos et al. 2024a With LoRA incorporated, J-LRD continues to show minimal performance degradation with an average drop of 1.00%. G-LRD (gs=4) and M-LRD show improved results compared to their non-LoRA counterparts, with average drops of 2.01% and 5.14%, respectively. Notably, with LoRA integration, G-LRD shows only a 1.03% accuracy difference compared to J-LRD.
Appendix G More Results on Zero-shot Accuracy
Following up Sec. 4.1, we further report the zero-shot evaluation results of Palu on the LLama2 family at 30% compression rate. As shown in Table. 7, we observe that Palu achieve nearly no accuracy drops using either J-LRD, G-LRD, or M-LRD. Thus, the users may adopt M-LRD first to optimize the efficiency further.
Table 8: Ablation study of low-rank decomposition group size on perplexity for the Llama2-7B model at a 50% compression rate using Wikitext-2.
Method Group Size Perplexity
Baseline - 5.47
J-LRD 32 5.62
G-LRD 16 5.74
8 5.88
4 6.01
2 6.42
M-LRD 1 6.81
Appendix H Ablation Study
H.1 Influence of Different Group Size
Since our proposed G-LRD method allows for balancing performance and efficiency by adjusting the group size, we conducted an ablation study on group size. As seen in Table 8, as the group size increases, the amount of shared information also increases, leading to improved performance.
H.2 Influence of Walsh-Hadamard Transform
We conduct the ablation study to profile the benefits of applying the Walsh-Hadamard Transform (WHT). Experiment results are reported at Table 9. On the 3-bit quantization level, we observe that the Hadamard Transform only brings a slight amount of perplexity. However, when we quantize the low-rank representation more extremely (i.e., 2-bit), we can observe a notable 4.17 perplexity enhancements. It’s worth re-emphasizing that Hadamard Transform will not bring extra overhead during inference, as Palu optimizes the WHT process via offline preprocessing. The reader may refer to Sec. 3.4 for more details.
H.3 Automatic Rank Allocation vs. Uniform Rank Allocation
Table 10 presents the ablation study on the impact of different rank allocation schemes on the model’s accuracy. Applying rank searching results in a notable performance improvement. For instance, at a compression rate of 50%, there is a significant reduction in perplexity by 2.18. Fig. 7 visualizes the rank allocation across different transformer blocks for key and value projection layers. The results clearly demonstrate a non-uniform allocation result. Specifically, we observe that the value is generally allocated a higher rank compared to the key. Additionally, the first half of the layers are assigned higher ranks, indicating their greater importance in preserving model performance. This visualization underscores the effectiveness of our rank search algorithm in identifying and allocating appropriate ranks to different components, thereby optimizing the balance between compression and accuracy.
Table 9: Ablation Study on different quantization settings for quantizing low-rank latent representations. Same as Sec. 4.2, we use the WikiText-2 with sequence length set to 4096 as the evaluation benchmark.
Method Wikitext-2 PPL $\downarrow$
Llama2-7B 5.12
Palu-30% (FP16) 5.25
      + 3-bits w/o Hadamard 5.52
      + 3-bits w Hadamard 5.33 (0.19$\downarrow$)
      + 2-bits w/o Hadamard 9.48
      + 2-bits w Hadamard 5.77 (3.71$\downarrow$)
Palu-50% (FP16) 5.63
      + 3-bits w/o Hadamard 5.99
      + 3-bits w Hadamard 5.77 (0.22$\downarrow$)
      + 2-bits w/o Hadamard 10.58
      + 2-bits w Hadamard 6.41 (4.17$\downarrow$)
Figure 7: Visualization of layer-wise low-rank compression rate on Llama2-7B with 50% of overall compression rate. Here, compression rates (i.e., rank) are allocated using the proposed Fisher Information-based automated rank allocation algorithm.
Table 10: Ablation study on w/ and w/o rank search. We use Llama2-7b and Wikitext-2 with sequence length 2048 as the benchmark.
Rate=30% Rate=50% Rate=70%
Uniform 6.34 7.36 10.77
Automatic (ours) 5.62 (0.72$\downarrow$) 6.02 (1.36$\downarrow$) 8.59 (2.18$\downarrow$)
Appendix I Experiment Details
I.1 Zero-Shot Evaluation Details
We selected six zero-shot tasks from the LM-eval benchmark to evaluate Palu:
We report accuracy for WinoGrande, PIQA, and ARC-Easy, and accuracy normalized by sequence length (acc_norm) for HellaSwag and ARC-Challenge.
I.2 LongBench Evaluation Details
For the LongBench evaluation in the manuscript, we selected eight tasks from four subgroups, ensuring a comprehensive evaluation of Palu. The tasks and their corresponding metrics are detailed below:
During evaluation, we set the maximum sequence length to 8192 for the Mistral model and 15500 for Vicuna.

### v2

@@
-Palu: Compressing KV-Cache with Low-Rank Projection
+Palu: KV-Cache Compression with Low-Rank Projection
 Abstract
-KV-Cache compression methods generally sample a KV-Cache of effectual tokens or quantize it into lower bits.
-However, these methods cannot exploit the redundancy of the hidden dimension of KV tensors.
-This paper investigates a unique hidden dimension approach called Palu, a novel KV-Cache compression framework that utilizes low-rank projection.
-Palu decomposes the linear layers into low-rank matrices, caches the smaller intermediate states, and reconstructs the full keys and values on the fly.
-To improve accuracy, compression rate, and efficiency, Palu further encompasses (1) a medium-grained low-rank decomposition scheme, (2) an efficient rank search algorithm, (3) a low-rank-aware quantization algorithm, and (4) matrix fusion with optimized GPU kernels.
-Our extensive experiments with popular LLMs show that Palu can compress KV-Cache by more than 91.25% while maintaining a significantly better accuracy (up to 1.19 lower perplexity) than state-of-the-art KV-Cache quantization methods at a similar or even higher memory usage.
-When compressing KV-Cache for 50%, Palu delivers up to 1.61$\times$ end-to-end speedup for the attention module.
-Our code is publicly available at: https://github.com/shadowpa0327/Palu.
-Palu: Compressing KV-Cache with Low-Rank Projection
-[…]
-footnotetext: These authors contributed equally to this work.
+Post-training KV-Cache compression methods typically either sample a subset of effectual tokens or quantize the data into lower numerical bit width.
+However, these methods cannot exploit redundancy in the hidden dimension of the KV tensors.
+This paper presents a hidden dimension compression approach called Palu, a KV-Cache compression framework that utilizes low-rank projection to reduce inference-time LLM memory usage.
+Palu decomposes the linear layers into low-rank matrices, caches compressed intermediate states, and reconstructs the full keys and values on the fly.
+To improve accuracy, compression rate, and efficiency, Palu further encompasses (1) a medium-grained low-rank decomposition scheme, (2) an efficient rank search algorithm, (3) low-rank-aware quantization compatibility enhancements, and (4) optimized GPU kernels with operators fusion.
+Extensive experiments with popular LLMs show that Palu compresses KV-Cache by 50%, while maintaining strong accuracy and delivering up to 1.89$\times$ speedup on the RoPE-based attention module.
+When combined with quantization, Palu’s inherent quantization-friendly design yields small to negligible extra accuracy degradation, while saving additional memory than quantization-only methods and achieving up to 2.91$\times$ speedup for the RoPE-based attention.
+Moreover, it maintains comparable or even better accuracy (up to 1.19 lower perplexity) compared to quantization-only methods.
+These results demonstrate Palu’s superior capability to effectively address the efficiency and memory challenges of LLM inference posed by KV-Cache.
+Our code is publicly available at: https://github.com/shadowpa0327/Palu
+$\ast$\$$\ast$\$footnotetext: Equal contribution\$$\dagger$\$$\dagger$\$footnotetext: Correspondence to: Chi-Chih Chang, cc2869@cornell.edu
 1 Introduction
-Large language models (LLMs) are revolutionizing the AI industry and providing a high-level intelligence that previous machine learning (ML) models could not achieve.
-To speed inference, caching key-value states (KV-Cache) in memory is a simple yet effective technique.
-However, the size of the KV-Cache can grow rapidly, straining memory capacity and bandwidth; further, the memory-bounded characteristics of the decoding stage also limit inference speed when loading KV-Cache data.
-Therefore, KV-Cache compression has become a central research topic for building efficient LLMs.
+Large language models (LLMs) have propelled AI into new applications and capabilities, providing a high-level intelligence that previous machine learning (ML) models could not achieve.
+To speed up inference, caching the Key-Value states (KV-Cache) in memory is a simple yet effective technique.
+However, the size of the KV-Cache can grow rapidly, straining memory capacity and bandwidth especially with long context lengths (Fu, 2024); further, the memory-bounded nature of the decoding stage limits inference speed when loading KV-Cache data (Gholami et al., 2024).
+Therefore, KV-Cache compression has become a central research topic for running LLMs efficiently.
+Although emerging attention mechanisms such as Multi-Query Attention (MQA) (Shazeer, 2019), Group-Query Attention (GQA) (Ainslie et al., 2023) and Multi-head Latent Attention (MLA) (DeepSeek-AI et al., 2024) can reduce KV-Cache size, it either requires model pre-training or has a significant impact on model’s accuracy when converting from traditional Multi-Head Attention (MHA) (Chen et al., 2024).
+In contrast, post-training KV-Cache compression techniques offer an alternative approach to advance efficiency for existing models.
+Among various KV-Cache compression methods, quantization (Liu et al., 2024b; Hooper et al., 2024) and token eviction (Zhang et al., 2024; Xiao et al., 2024) stand out as effective strategies to reduce the memory footprint of KV-Cache.
+Quantization methods aim to reduce the bit-width used to represent each piece of data, while token eviction techniques focus on retaining a partial set of KV-Cache.
+However, both methods neglect the hidden dimensions of the KV-Cache, where substantial redundancy often resides.
+To capitalize on this untapped potential, we introduce Palu, a post-training KV-Cache compression framework that leverages low-rank projection to reduce the hidden dimension of KV tensors, offering an additional and orthogonal compression dimension to existing quantization and token eviction methods.
+A naive way to utilize low-rank projection for compressing the KV-Cache is by directly mapping cached matrices into low-rank space (Jolliffe & Cadima, 2016; Zhao et al., 2024).
+However, this approach imposes an unacceptably heavy overhead of computing the decomposition matrices during runtime.
+To avoid this, Palu statically decomposes the Key and Value-projection weight matrices and caches the latent representations of the low-rank decomposition (see Fig.
+1).
+This innovative design enables Palu to reduce memory while mitigating the runtime overhead of KV-Cache low-rank decomposition.
 Figure 1: Palu’s low-rank projection method for KV-Cache reduction.
@@
 Input X is down-projected to latent representation H, which is cached.
-Y is reconstructed from H on-the-fly using the up-projection matrix B.
-Research on KV-Cache compression falls into two categories: quantization and token eviction.
-Quantization methods aim to use less bit-width to represent each datum; token eviction techniques seek ways to retain only a partial set of KV-Cache.
-However, both categories fail to explore the hidden dimensions of KV-Cache where high redundancy often occurs.
-This work examines the problem from a fundamentally different perspective, i.e., ways to reduce the hidden dimensions.
-We introduce Palu, a KV-Cache compression framework that reduces the size of the KV-Cache through low-rank projection.
-Although low-rank projection has been shown to improve LLM efficiency, e.g., by reducing model size (Yuan et al., 2023; Wang et al., 2024) or minimizing memory footprint for finetuning (Hu et al., 2022; Dettmers et al., 2023), it has not been studied for compressing the KV-Cache, especially in post-training.
-A naive way to compress the KV-Cache with low-rank projection is by directly mapping cached matrices into low-rank space (Jolliffe and Cadima, 2016; Zhao et al., 2024).
-However, this approach imposes an unacceptably heavy overhead of computing the decomposition matrices during runtime, after the prefill stage.
-To avoid this, Palu statically decomposes the key and value-projection weight matrices and caches the latent representations of the low-rank decomposition (Fig.
-1), instead of directly decomposing the KV-Cache.
-This innovative design enables Palu to reduce memory while substantially mitigating runtime overhead.
-Despite its promising memory reduction, we found that naively decomposing the key and value projection matrices along the head dimension causes a substantial loss in accuracy.
-Although we found that decomposing the entire matrix at one time can maintain accuracy, it significantly increases reconstruction overhead.
-To find a balance, Palu proposes a medium-grained, group-head low-rank decomposition optimization that is both accurate and less costly in terms of reconstruction.
-For LLMs, each linear projection module has a different sensitivity to compression Sharma et al.
-(2023); Yuan et al.
-(2023).
-To exploit the sensitivity and improve accuracy, we design an efficient rank search algorithm based on Fisher information.
-Our algorithm automatically assigns a larger rank for important matrices and smaller ranks for less critical ones, boosting accuracy at the same compression rate.
-In addition to low-rank decomposition, we integrate quantization into Palu.
-We found that low-rank decomposition often induces severe outliers in the latent representation.
-Inspired by recent LLM quantization methods Tseng et al.
-(2024); Ashkboos et al.
-(2024b), we use Hadamard transformation matrices with a low-rank-aware quantization algorithm to eliminate outliers and increase quantization accuracy.
-To avoid runtime overhead, Palu fuses the transformation matrices into the forward and backward decomposed matrices, concealing the overhead completely.
-To enhance inference latency, we integrate the decomposed matrix into the subsequent linear layer, thereby mitigating the online reconstruction overhead.
-In scenarios where RoPE is applied and matrix fusion is infeasible, we developed an optimized GPU kernel to accelerate the reconstruction.
-We evaluate Palu on popular LLMs and benchmarks.
-Our experiments show that Palu maintains strong zero-shot accuracy and perplexity with up to 50% low-rank compression.
-Combining low-rank compression and quantization, Palu compresses KV-Cache by over 91.25% (11.4$\times$ compression) while maintaining perplexity.
-Compared to a state-of-the-art KV-Cache quantization (Hooper et al., 2024) with 87.5% compression, Palu achieves a significantly lower perplexity of 1.19.
-Additionally, with 50% compression without quantization, Palu demonstrates a 1.61$\times$ end-to-end speedup, highlighting Palu’s ability to reduce memory usage and enhance speed for LLMs.
+Y can be reconstructed from H using the up-projection matrix B.
+In designing an effective decomposition strategy for attention modules with multiple attention heads, we observed a clear trade-off between accuracy and reconstruction overhead.
+Decomposing the projection matrices across all attention heads together improves accuracy by preserving global information, but this approach significantly increases reconstruction costs.
+On the other hand, decomposing each head separately reduces reconstruction overhead but leads to a higher loss in accuracy.
+To address this, Palu introduces a medium-grained, group-head low-rank decomposition that strikes a balance between accuracy and reconstruction efficiency.
+For LLMs, each linear projection module has a different sensitivity to compression (Sharma et al., 2023; Yuan et al., 2023).
+To exploit the sensitivity and improve accuracy, we design an efficient rank search algorithm based on Fisher information (Ly et al., 2017; Liu et al., 2021).
+Our algorithm automatically assigns a higher rank for important matrices and lower ranks for less critical ones, boosting accuracy at the same overall KV-Cache compression rate.
+In addition to its low-rank decomposition, Palu is compatible with quantization techniques.
+We found that low-rank decomposition can introduce severe outliers in the latent representation, which significantly hinders accurate low-bit quantization.
+Although the Hadamard transformation has been shown to be effective for outlier elimination in recent studies (Tseng et al., 2024; Ashkboos et al., 2024b; Liu et al., 2024a; Chiang et al., 2024), its integration often introduces computational overhead during runtime.
+However, Palu’s inherent matrix pair structure makes it highly compatible with this technique, allowing the transformation matrices to be seamlessly fused into the forward and backward matrices, effectively mitigating outliers without impacting runtime efficiency.
+We evaluate Palu on widely used LLMs and benchmarks.
+Our experiments demonstrate that Palu maintains strong zero-shot accuracy and perplexity with up to 50% low-rank compression.
+Moreover, when combining low-rank compression with quantization, Palu achieves an impressive over 91.25% compression (11.4$\times$ reduction) and yields a significantly lower perplexity of 1.19 than KVQuant (Hooper et al., 2024), a state-of-the-art KV-Cache quantization method, which only achieves an 87.5% compression rate.
+For latency evaluation, under a 50% KV-Cache compression rate without quantization, Palu demonstrates up to 1.89$\times$ and 2.2$\times$ speedup for RoPE-based and non-RoPE attention modules.
+When integrated with quantization, Palu achieves up to 2.91$\times$ and 6.17$\times$ acceleration on RoPE-based and non-RoPE attention, respectively.
+These results underscore Palu’s ability to significantly reduce KV-Cache memory footprint while boosting inference efficiency for LLMs.
 Our key contributions include:
-  • Palu, a new post-training KV-Cache compression framework that caches low-rank latent representations.
-  • Group-head low-rank decomposition (G-LRD), an optimization for balancing accuracy and efficiency.
+  • Palu, a new post-training KV-Cache compression framework that caches low-rank latent representations of Key and Value states.
+  • Group-head low-rank decomposition (G-LRD), an optimization for balancing accuracy and reconstruction efficiency.
   • An automated rank search algorithm for adaptively assigning ranks to each decomposed matrix, given a target compression rate.
-  • A low-rank-aware quantization algorithm that eliminates low-rank-induced outliers and imposes zero runtime overhead.
+  • A co-designed quantization compatibility optimization that eliminates low-rank-induced outliers and imposes zero runtime overhead.
 2 Background
@@
 2.2 Singular Value Decomposition (SVD)
-SVD (Jolliffe and Cadima, 2016) is a commonly used technique for computing the low-rank approximation for a given matrix.
+SVD (Jolliffe & Cadima, 2016) is a commonly used technique for computing the low-rank approximation for a given matrix.
 We now introduce our specific use of SVD for Palu’s default low-rank decomposition method.
-Given a weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, SVD decomposes $\mathbf{W}$ into three matrices: $\mathbf{W}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{T}$.
+Given a weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$, SVD decomposes $\mathbf{W}$ into three matrices: $\mathbf{W}=\mathbf{U}\bm{\Sigma}\mathbf{V}^{T}$.
 Here, $\mathbf{U}$ and $\mathbf{V}$ are orthogonal matrices containing the left and right singular vectors, respectively.
-The matrix $\boldsymbol{\Sigma}$ is a diagonal matrix that consists of singular values.
+The matrix $\bm{\Sigma}$ is a diagonal matrix that consists of singular values.
 We describe the decomposition as
-$\mathbf{W}\approx\mathbf{A}\mathbf{B},\hskip 5.0pt\mathbf{A}=\mathbf{U}_{r}\sqrt{\boldsymbol{\Sigma}_{r}},\hskip 5.0pt\mathbf{B}=\sqrt{\boldsymbol{\Sigma}_{r}}\mathbf{V}_{r}^{T},$
-where $\mathbf{A}\in\mathbb{R}^{m\times r}$, $\mathbf{B}\in\mathbb{R}^{r\times n}$, $\boldsymbol{\Sigma}_{r}\in\mathbb{R}^{r\times r}$.
-$\boldsymbol{\Sigma}_{r}$ is a diagonal matrix containing the largest $r$ singular values, and $\mathbf{U}_{r}$, $\mathbf{V}_{r}^{T}$ are corresponding singular vectors truncated from $\mathbf{U}$ and $\mathbf{V}^{T}$.
+$\mathbf{W}\approx\mathbf{A}\mathbf{B},\hskip 5.0pt\mathbf{A}=\mathbf{U}_{r}\sqrt{\bm{\Sigma}_{r}},\hskip 5.0pt\mathbf{B}=\sqrt{\bm{\Sigma}_{r}}\mathbf{V}_{r}^{T},$
+where $\mathbf{A}\in\mathbb{R}^{m\times r}$, $\mathbf{B}\in\mathbb{R}^{r\times n}$, $\bm{\Sigma}_{r}\in\mathbb{R}^{r\times r}$.
+$\bm{\Sigma}_{r}$ is a diagonal matrix containing the largest $r$ singular values, and $\mathbf{U}_{r}$, $\mathbf{V}_{r}^{T}$ are corresponding singular vectors truncated from $\mathbf{U}$ and $\mathbf{V}^{T}$.
 This truncation and subsequent matrix formation let us approximate weight matrix $\mathbf{W}$ with two low-rank matrices $\mathbf{A}$ and $\mathbf{B}$, thereby reducing the storage by $\frac{mr+rn}{mn}$.
@@
 3.1 Compressing the KV-Cache via Low-Rank Projection
-As noted, a naive approach to compress the KV-Cache with low-rank projection is to apply SVD directly on the KV-Cache and store the top-$r$ singular vectors.
-However, this approach poses significant computational challenges during runtime that make it impractical for deployments (see Section 1).
-To apply low-rank projection more efficiently, Palu uses SVD to decompose the key and value projection weight matrices.
+To apply low-rank projection more efficiently than directly decomposing the KV-Cache during runtime, Palu uses SVD to decompose the Key and Value projection matrices.
 This approach is based on the observation that low-rank decomposition rewrites the linear projection layer from $\mathbf{y}=\mathbf{x}\mathbf{W}$ into $\mathbf{y}=\mathbf{x}\mathbf{A}\mathbf{B}$.
@@
 $\mathbf{h}=\mathbf{A}\mathbf{x},\quad\mathbf{y}=\mathbf{B}\mathbf{h}$ (4)
-This two-step process lets Palu (1) store the lower dimension latent representation instead of the origin keys and values states, and (2) reconstruct them during decoding.
+This two-step process lets Palu (1) store the lower dimension latent representation instead of the origin key and value states, and (2) reconstruct them during decoding.
 3.1.1 Integration with the Attention Mechanism and Offline Matrix Fusion
@@
 For each attention head $i$, Palu applies SVD and maps the key-projection matrix $\mathbf{W}^{k}_{i}$ and value-projection matrix $\mathbf{W}^{v}_{i}$ into $\mathbf{A}^{k}_{i}\mathbf{B}^{k}_{i}$ and $\mathbf{A}^{v}_{i}\mathbf{B}^{v}_{i}$.
-Based on the formula of attention output in equation 2, Palu absorbs the reconstruction matrix $\mathbf{B}^{v}_{i}$ into the output projection matrix $\mathbf{W}_{i}^{o}$ offline:
-$\mathbf{a}_{i}\mathbf{W}^{o}_{i}$ $=(\mathbf{p}_{i}\mathbf{V}_{i})\mathbf{W}^{o}_{i}$ (5)
-$=(\mathbf{p}_{i}\mathbf{H}^{v}_{i}\mathbf{B}^{v}_{i})\mathbf{W}^{o}_{i}$
-$=\mathbf{p}_{i}\mathbf{H}^{v}_{i}(\mathbf{B}^{v}_{i}\mathbf{W}^{o}_{i})$
+Based on the formula of attention output in Eq. 2, Palu absorbs the reconstruction matrix $\mathbf{B}^{v}_{i}$ into the output projection matrix $\mathbf{W}_{i}^{o}$ offline:
+$\mathbf{a}_{i}\mathbf{W}^{o}_{i}$ $=(\mathbf{p}_{i}\mathbf{V}_{i})\mathbf{W}^{o}_{i}=(\mathbf{p}_{i}\mathbf{H}^{v}_{i}\mathbf{B}^{v}_{i})\mathbf{W}^{o}_{i}=\mathbf{p}_{i}\mathbf{H}^{v}_{i}(\mathbf{B}^{v}_{i}\mathbf{W}^{o}_{i})$ (5)
 Such fusion lets Palu skip the explicit reconstruction of the full value vectors, reduce the number of matrix multiplications, and improve efficiency.
@@
 Matrix $\mathbf{B}^{k}_{i}$ can be fused into the query projection matrix $\mathbf{W}_{i}^{q}$ offline, as shown by
-$\mathbf{q}_{i}\mathbf{K}_{i}^{T}$ $=\mathbf{q}_{i}(\mathbf{H}^{k}_{i}\mathbf{B}^{k}_{i})^{T}$ (6)
-$=\mathbf{x}_{t}\mathbf{W}^{q}_{i}(\mathbf{B}^{k}_{i})^{T}(\mathbf{H}^{k}_{i})^{T}$
-$=\mathbf{x}_{t}\Bigl{(}\mathbf{W}^{q}_{i}(\mathbf{B}^{k}_{i})^{T}\Bigr{)}(\mathbf{H}^{k}_{i})^{T}.$
+$\mathbf{q}_{i}\mathbf{K}_{i}^{T}$ $=\mathbf{q}_{i}(\mathbf{H}^{k}_{i}\mathbf{B}^{k}_{i})^{T}=\mathbf{x}_{t}\mathbf{W}^{q}_{i}(\mathbf{B}^{k}_{i})^{T}(\mathbf{H}^{k}_{i})^{T}=\mathbf{x}_{t}\Bigl{(}\mathbf{W}^{q}_{i}(\mathbf{B}^{k}_{i})^{T}\Bigr{)}(\mathbf{H}^{k}_{i})^{T}.$ (6)
 Here, $\mathbf{B}^{k}_{i}\in\mathbb{R}^{r\times d_{h}}$ and $\mathbf{W}^{q}_{i}\in\mathbb{R}^{d\times d_{h}}$, so the fused matrix $(\mathbf{W}^{q}_{i}(\mathbf{B}^{k}_{i})^{T})$ has size $\mathbb{R}^{d\times r}$.
 This fusion boosts computational efficiency by reducing the matrix dimension during attention score calculation.
-3.1.2 Mechanism with Positional Embedding
-Recent LLMs, such as the Llama family, apply positional embedding (i.e., RoPE (Su et al., 2021)) onto the query and key states before multiplication.
-The non-linear operations of adding positional embedding prevent the fusion for the attention score, as described in equation 6.
-In this case, Palu reconstructs the keys from latent representations on-the-fly during decoding.
-We improve reconstruction efficiency by designing an optimized kernel, which we present in Sec.
-3.5.
-Note that for some positional embedding methods, such as ALiBi Press et al.
-(2022), or new attention mechanisms, such as MLA DeepSeek-AI et al.
-(2024), positional embedding is not directly applied on the key states; therefore, the fusion of equation 6 remains valid.
-We expect the end-to-end speedup to be larger on these models.
-Fig.
-2 presents a detailed workflow of Palu based on Llama’s attention with RoPE.
+3.1.2 Compability with Positional Embedding
+Recent LLMs, such as the Llama family, apply Rotary Positional Embedding (i.e., RoPE (Su et al., 2021)) onto the Query and Key states prior to their multiplication.
+The non-linear nature of these positional embeddings prevents the matrix fusion of attention scores, as outlined in Eq.
+6.
+To address this, Palu dynamically reconstructs the keys from latent representations during decoding.
+We enhance reconstruction efficiency with a custom-designed GPU kernel, detailed in Sec.
+4.1 and further described in Appendix B.
+Note that for some positional embedding methods, such as ALiBi (Press et al., 2022), or new attention mechanisms, such as MLA (DeepSeek-AI et al., 2024), positional embedding is not directly applied to the Key states.
+Consequently, the fusion described in Eq.
+6 remains valid.
+For these non-RoPE attention modules, Palu achieves greater speedup compared to RoPE-based attention, as their reconstruction can be avoided with matrix fusion.
+A detailed workflow of Palu in the context of Llama’s RoPE-based attention is illustrated in Fig.
+2.
 3.2 Decomposition Granularity
 3.2.1 Multi-Head Low-Rank Decomposition
-We name the per-head decomposition scheme in Sec.3.1.1 as multi-head low-rank decomposition (M-LRD).
-We found M-LRD often causes a non-negligible accuracy degradation (discussed further in Sec.4.1), possibly because SVD fails to capture the common information shared across heads.
-Therefore, alternative approaches are needed for preserving model accuracy.
+We name the per-head decomposition scheme in Sec.
+3.1.1 as multi-head low-rank decomposition (M-LRD).
+We found M-LRD often causes a non-negligible accuracy degradation (discussed further in Sec.
+4.2), possibly because SVD fails to capture the common information shared across heads.
+Therefore, alternative approaches are needed to preserve model accuracy.
 3.2.2 Joint-Head Low-Rank Decomposition
@@
 High Inference Overhead.
-Despite better preserving model information, J-LRD introduces significant computational and memory overhead during decoding.
-Specifically, the total number of floating point operations (FLOPs) to reconstruct the key or values state of one head now becomes $r_{\text{joint}}\cdot d_{h}\cdot n$.
+Despite better preserving model accuracy, J-LRD introduces significant computational and memory overhead during decoding.
+Specifically, the total number of floating point operations (FLOPs) to reconstruct the Key or Value state of one head now becomes $r_{\text{joint}}\cdot d_{h}\cdot n$.
 Assuming the same size as the total low-rank latent representations (i.e., $r_{\text{joint}}=\sum_{i=1}^{n}{r_{i}}$), the total reconstruction cost is $n$ times higher than M-LRD, whose total FLOPs is $r_{i}\cdot d_{h}\cdot n$.
-When considering the matrix fusion in Sec.3.1.1, the fused matrix size of J-LRD has a size of $r_{joint}\cdot d\cdot n$, which is also $n$ times larger than M-LRD, leading to a substantial higher memory consumption.
+When considering the matrix fusion in Sec.
+3.1.1, the fused matrix of J-LRD has a size of $r_{joint}\cdot d\cdot n$, which is also $n$ times larger than M-LRD, leading to substantial higher memory consumption.
 Figure 3: Performing decomposition at different granularities.
@@
 To sum up, G-LRD offers a middle ground between computation overhead and approximation accuracy.
-We illustrate M-LRD, J-LRD and G-LRD in Fig.3.
-For further discussion on decomposed weight matrix size, please refer to Appendix B, where more detailed information is provided.
+We illustrate M-LRD, J-LRD and G-LRD in Fig.
+3.
+Please refer to Appendix C for further discussions on the costs of different decomposition granularities.
 3.3 Automatic Rank Allocation
@@
 In Palu, we identify Fisher information (Ly et al., 2017; Liu et al., 2021) as an accurate approximator since it can quantify the amount of information for each parameter.
-We then employ the sum of Fisher information to estimate the importance of the weight matrix of each linear layer.
+We then employ the sum of Fisher information to estimate the importance of the weight matrix of each linear layer (Abdelfattah et al., 2021).
 Assuming that the compression sensitivity is proportional to Fisher information, we determine the rank for each weight matrix by computing the ratio of its Fisher information to the total Fisher information across all decomposition targets.
 We use this ratio to allocate the compression rate (i.e., rank level $r$), ensuring that more important layers retain higher rank levels.
-For a detailed ablation study on our Automatic Rank Allocation method, please refer to Appendix H.3, where we analyze the impact of our strategy on model performance.
-3.4 Low-Rank-Aware Quantization
-Figure 4: Activation distribution of the low-rank compressed key caches at the $4^{th}$ Llama-2 attention layer.
-We integrate quantization into Palu to further compress the KV-Cache.
-However, we observe that low-rank compressed latent representations have severe outliers, which limit quantization applicability in Palu.
-Unlike natural outliers described in previous KV-Cache quantization literature (Liu et al., 2024; Hooper et al., 2024), these outliers are induced by SVD-based low-rank factorization.
+For a detailed ablation study on our Automatic Rank Allocation, please refer to Appendix G.3.
+3.4 Quantization Compatibility
+Figure 4: Activation distribution of the low-rank key caches at the $4^{th}$ Llama-2 attention layer.
+We integrate quantization into Palu to compress the KV-Cache further.
+We observe that low-rank compressed latent representations have severe outliers, which limit quantization applicability in Palu.
+Unlike natural outliers described in previous KV-Cache quantization literature (Liu et al., 2024b; Hooper et al., 2024), these outliers are induced by SVD-based low-rank factorization.
 Fig.
@@
 This pattern stretches the data distribution and hurts quantization accuracy.
-Inspired by recent LLM quantization literature (Ashkboos et al., 2024b; Tseng et al., 2024), we apply the Walsh-Hadamard transform (WHT, Fino and Algazi) to eliminate outliers (Fig.
+Inspired by recent LLM quantization literature (Ashkboos et al., 2024b; Tseng et al., 2024), we apply the Walsh-Hadamard transform (WHT, Fino & Algazi) to eliminate outliers (Fig.
 4 (b)), enabling a high quantization accuracy.
 However, this transformation introduces an extra matrix multiplication with associated runtime overhead.
-Unlike earlier methods (Ashkboos et al., 2024b) that must apply online WHT when quantizing KV-Cache, we optimize this process by integrating the Hadamard matrix into low-rank decomposed weights with no additional overhead, as described by
+Unlike earlier methods (Ashkboos et al., 2024b) that must apply online WHT when quantizing KV-Cache, we optimize this process by integrating the Hadamard matrix into low-rank decomposed weights with no additional compute overhead, as described by
 $\mathbf{W}\approx\mathbf{AB}=(\mathbf{AR})(\mathbf{R}^{T}\mathbf{B})=\mathbf{\hat{A}\hat{B}},$ (7)
 where $\mathbf{R}$ is the Hadamard matrix.
-This optimization allows Palu can integrate the proposed low-rank compression technique with low-bit quantization.
-Our experiments show that, on top of the low-rank compression, our quantization method only negligibly increases perplexity, even at extreme levels such as 3-bit or 2-bit (see Sec.4.2).
-3.5 Efficient Inference Implementation
-Support for RoPE-based Attention.
-To support RoPE-based attention, we fuse the reconstruction of original keys, applying positional embedding (i.e., RoPE (Su et al., 2024)) and the final matrix-vector multiplication of query and reconstructed keys by implementing customized GPU kernels in Triton Tillet et al.
-(2019).
-We leverage the tiling techniques akin to FlashAttention (Dao et al., 2022) to minimize the memory reads/writes of GPU high bandwidth memory.
-Table 1: Perplexity and zero-shot accuracy of Palu, with different decomposition strategies at 50% compression rate.
+This optimization allows Palu to integrate the proposed low-rank compression technique with low-bit quantization.
+Our experiments show that, on top of the low-rank compression, our quantization method only negligibly increases perplexity, even at extreme levels such as 3-bit or 2-bit with a simple per-token quantization scheme (see Sec.
+4.3).
+4 Experiments
+4.1 Experiments Setup
+Models and Tasks.
+We evaluate Palu on four LLM families, Llama-2 (Touvron et al., 2023), Llama-3 (Dubey et al., 2024), Mistral (Jiang et al., 2023) and LongChat (Li et al., 2023).
+For accuracy evaluation, we measure perplexity on the WikiText-2 (Merity et al., 2016) and C4 (Raffel et al., 2020) datasets and use LM-Evaluation-Harness (Gao et al., 2023) to measure zero-shot accuracy on six common sense tasks.
+We also evaluate long context accuracy on 16 tasks in LongBench (Bai et al., 2023).
+Unless specification, we refer to baseline as a model with non-compressed KV-Cache.
+See Appendix H for further details on the dataset and settings.
+Compression Settings.
+We implemented Palu based on the Huggingface library (Wolf et al., 2020).
+Decomposition of the Key and Value projection layers was performed using the truncation-aware SVD method proposed by SVD-LLM (Wang et al., 2024).
+Unless otherwise specified, Palu’s results are G-LRD with a group size of 4 (gs-4), with equal rank size for each group.
+To calculate Fisher information in rank searching, we used 2048 random samples from Wikitext-2, each with a sequence length of 1024.
+For quantization integration in Palu, we use a simple per-token, asymmetric integer quantization.
+For evaluation on quantization results, we compare Palu to advanced KV-Cache quantization methods, including Atom (Zhao et al., 2023), KVQaunt (Hooper et al., 2024), and KIVI (Liu et al., 2024b).
+Refer to Sec.
+5 for a brief summary of these methods.
+GPU Kernels Implementation.
+[…]
+4.2 Results with Different Decomposition Granularity
+We evaluate perplexity and zero-shot accuracy of Palu with a 50% low-rank compression rate using M-LRD, G-LRD, and J-LRD on Llama-2-7B and Llama-3-8B-Instruct, and present the results in Table 1.
+Table 1: Perplexity and zero-shot accuracy of Palu at 50% compression rate.
 Model Method Perplexity $\downarrow$ Zero-Shot Accuracy (%) $\uparrow$
 Wiki2 C4 OBQA Hella PIQA ARC-e ARC-c Wino Avg.
-Llama2-7B Baseline 5.47 7.26 44.20 76.00 78.07 76.30 46.42 69.30 65.05
+Llama-2-7B Baseline 5.47 7.26 44.20 76.00 78.07 76.30 46.42 69.30 65.05
 J-LRD 5.62 7.75 45.40 75.57 77.48 75.97 45.31 69.22 64.82
@@
 M-LRD 6.75 12.01 39.60 65.35 74.76 67.17 35.24 64.64 57.79
-Llama2-13B Baseline 4.88 6.70 45.20 79.39 79.11 79.42 49.06 72.38 67.43
-J-LRD 4.97 6.92 46.40 79.48 78.62 79.29 49.91 70.56 67.38
-G-LRD 5.31 7.76 45.60 77.29 77.42 76.05 45.99 72.45 65.80
-M-LRD 5.65 8.34 43.20 74.34 77.53 75.76 45.39 68.98 64.20
-4 Experiments
-Table 2: Accuracy of Palu with 30% and 50% compression rate on 8 tasks in LongBench.
-Model Method SAMSum LCC M-News Qasper TREC QmSum R-Bench TriQA Avg.
-Vicuna-7B Baseline 40.73 49.20 27.98 25.57 69.00 22.54 43.50 86.42 45.62
-Palu-30% 40.42 49.00 26.96 18.54 65.50 21.78 45.77 85.22 44.15
-Palu-50% 38.36 34.59 24.78 19.77 63.50 21.44 35.42 79.74 39.70
-Mistral-7B Baseline 41.71 66.59 20.26 8.13 67.00 19.99 59.10 89.90 46.59
-Palu-30% 43.23 64.23 23.50 10.21 72.50 21.16 56.80 89.82 47.68
-Palu-50% 36.43 50.22 19.23 7.54 67.50 19.56 44.76 85.37 41.33
-Models and Tasks.
-We evaluate Palu on three LLM families, Llama2(Touvron et al., 2023), Mistral(Jiang et al., 2023) and Vicuna(Chiang et al., 2023).
-For accuracy evaluation, we measure perplexity on the WikiText-2 (Merity et al., 2016) and C4 (Raffel et al., 2020) datasets and use LM-Evaluation-Harness (Gao et al., 2023) to measure zero-shot accuracy on six common sense tasks.
-We also evaluate long context accuracy on 8 tasks in LongBench (Bai et al., 2023).
-See Appendix I for further details on the dataset and settings.
-Compression Settings.
-We implemented Palu using the Huggingface library Wolf et al.
-(2020).
-Decomposition of the key and value projection layers was performed using the truncation-aware SVD method proposed by SVD-LLM (Wang et al., 2024).
-Unless otherwise specified, results for Palu are reported using G-LRD with a group size (gs) of 4 with equal rank size for each group.
-To calculate Fisher information in rank searching, we used 2048 samples from Wikitext-2, each with a sequence length of 1024.
-For quantization, we aligned our settings with Atom (Zhao et al., 2023), and use a simple per-token, asymmetric quantization to quantize the low-rank latent representation of both key and value states.
-4.1 Main Results
+Llama-3-8B-Inst Baseline 8.28 13.01 43.20 75.80 78.62 81.61 56.83 71.90 67.99
+J-LRD 9.12 15.90 43.40 73.20 76.50 79.63 51.96 72.45 66.19
+G-LRD 10.11 17.87 42.60 70.36 76.06 76.30 48.99 72.38 64.45
+M-LRD 12.38 23.02 38.80 63.04 73.67 69.78 42.58 62.51 58.40
 Perplexity Evaluation.
-We first present perplexity evaluation results on Wikitext-2 and C4 in Table 1.
-For the Llama2-7B model, we observe that the M-LRD method fails to maintain low perplexity at a 50% compression rate.
-In contrast, the J-LRD method performs significantly better given the same compression rate, achieving perplexity values of 5.62 on Wikitext-2.
-The G-LRD method, with a group size of 4, achieves an optimal balance between perplexity and computational cost, yielding perplexity values of 6.01 on Wikitext-2 at a 50% compression rate.
-The same trend is consistent across the Llama2-13B model.
+As Table 1 shows, for the Llama-2-7B model, Palu’s M-LRD method fails to maintain a low perplexity at a 50% compression rate.
+In contrast, despite having a high recomputation cost, J-LRD significantly outperforms M-LRD and achieves a 5.62 perplexity on WikiText-2.
+For G-LRD, which still maintains a low computation cost, yields a 6.01 perplexity on Wikitext-2, showing a great balance between model accuracy and compression overheads.
+The same trend is observed in the Llama-3-8B model as well.
+More results Llama-2-13B can be found in ppendix F.
 Zero-shot Evaluation Results.
-We further evaluate Palu on six common-sense zero-shot tasks, as shown in Table 1.
-Similar to the perplexity evaluation, the J-LRD method demonstrates the best performance on Llama2-7B, with only a 0.23% average accuracy degradation, while the M-LRD method results in the lowest average performance, with a 7.26% drop in accuracy compared to the baseline.
-In comparison, G-LRD experiences only a 2.4% average accuracy decline.
-LongBench Results.
-To assess whether Palu can effectively compress the KV-Cache for long-context scenarios, we present the performance of Palu across eight tasks from LongBench.
-We select Mistral-7B and Vicuna-7B (Chiang et al., 2023) as our evaluation targets, which support sequences up to 8k and 16k, respectively.
-Table 2 reports experiment results.
-We find that at a 50% compression level, it is relatively difficult to fully preserve accuracy.
-However, when compressing KV-Cache for 30%, Palu only has a 1.47% averaged accuracy degradation for Vicuna-7B, and even achieves a 1.09% enhancement for Mistral-7B.
-These results evidence of Palu’s ability to shrink the KV-Cache while maintaining performance.
-For detailed information on the integration of Palu with quantization and the performance measurements on LongBench, please refer to Appendix E.
-4.2 Integrating Palu with Quantization
-Table 3 shows the results of integrating quantization with Palu.
-With 3-bit quantization, Palu exhibits only a negligible perplexity increase of 0.08 at a 30% compression rate and a minimal 0.23 at a 50% compression rate, both compared to Palu at 16-bit.
-This indicates a minimal trade-off for substantial compression of integrating quantization.
-Furthermore, at 2-bit quantization, Palu significantly outperforms KVQuant, which is a SOTA quantization-based method.
-Palu demonstrates a compelling reduction in perplexity by 1.19 and 0.54 points, while consuming 30% and 50% less memory, respectively.
-These results underscore the effectiveness of Palu, establishing it as a superior KV-Cache compression strategy.
-4.3 Latency Evaluation
-We now provide the latency evaluation for Palu.
-In our experiment, latency is measured on an A6000 GPU, and the averaged results over 100 runs are reported.
-For baseline, we use the default implementation from HuggingFace in PyTorch.
-We do not apply FlashAttention (Dao et al., 2022) for an apple-to-apple comparison, as our current implementation does not fuse the Softmax operation.
-We leave the integration with FlashAttention as the future work.
-GEMV with Online Reconstruction.
-We evaluate the performance of our fused reconstruction kernel (Sec.
-3.5) for computing RoPE-based attention score.
-Specifically, we measure the latency from pre-RoPE query vector to the post-GEMV attention score (refer to Fig.
-2 for details).
-Here, we set the compression rate for our kernel as 75%.
-It’s because we observe the key states often achieve such compression rate when setting a 50% overall compression rate (see Appendix H.3 for detailed distribution).
-We compare Palu’s runtime against the baseline, which is simply a GEMV, and report the speedup in Fig.
-5 (a).
-For group size 32 (J-LRD), although it’s the most accurate decomposition, the high reconstruction leads to a significant 7.09$\times$ slowdown at sequence length 64K .
-For group size 1 (M-LRD), our fused kernel achieves up to a 2.48$\times$ speedup at a sequence length 16k, showcasing a high speedup when more accuracy loss is acceptable.
-For group size 2 and 4 (G-LRD), our fused kernel demonstrates 1.75$\times$ and 1.16$\times$ speedup at sequence length 16K.
-These results highlight the importance of exploring trade-off between decomposition granularity, reconstruction cost, and accuracy.
-Furthermore, we observe our kernel speedup decreases for sequence lengths over 16k, due to the increasing reconstruction cost shifts the attention operation from memory- to compute-bounded.
-Table 3: Perplexity and KV-Cache size comparison of Palu and other quantization-based methods for Llama2-7B on WikiText-2.
-For perplexity, sequence length is set to 4096, aligning with baselines.
-KV-Cache size is demonstrated for a sequence length of 128K.
+Similar to the perplexity evaluation, the J-LRD method demonstrates the best performance for the zero-shot accuracy on Llama-2-7B, with only a 0.23% average accuracy degradation.
+M-LRD method results in the lowest average performance, with a 7.26% drop in accuracy compared to the baseline.
+In comparison, G-LRD only has a 2.4% average accuracy decline, offering a sweet spot between model accuracy and compression overheads again.
+4.3 Results of Quantization Integration
+Table 2: Quantization perplexity and KV-Cache size for Llama-2-7B on WikiText-2. For perplexity, sequence length is 4096. KV-Cache size is demonstrated for 128K sequence length.
 Method Bit PPL
@@
 Palu-50% 2 6.41 4.0 93.75%
-End-to-end attention module.
-We measure the end-to-end latency of the entire RoPE-based attention module, with the integration of our fused reconstruction kernel for key states (Sec.
-3.5) and offline matrix fusion for value states (Eq.
-5).
-Here, we set the compression rate for key and value as 75% and 25%, for an overall 50% compression rate.
-The detailed forward pass operations are presented in Fig.
-2.
-We measure the decoding runtime for sequence length from 4K to 64K, and report the runtime of Palu and baseline in Fig.
-5 (b).
-As Fig.
-5 (b), Palu with smaller group sizes consistently delivers more than 1.43$\times$ speedup across all sequence lengths.
-For group size 1 (M-LRD), Palu achieves a 1.98$\times$, close to optimal (for 50% compression rate), speedup for sequence length 16K.
-For group size 4 (G-LRD), which is the sweet spot for maintaining strong accuracy and limiting reconstruction cost, Palu achieves up to 1.61$\times$ end-to-end speedup at sequence length 64K.
-We observe increasing speedup as sequence length increases from 4k to 64k due to continuously reducing latency from compressing value states.
-However, speedup is expected to converge or slightly decrease with further length increases due to higher reconstruction costs and the compute-bound nature of reconstruction kernels.
-We expect the speedup of Palu will be larger for non-RoPE-based attention, as the reconstruction overhead can be entirely eliminated by applying Eq.
-6.
-For the end-to-end model latency, please refer to Appendix D, where more detailed information and data are provided.
-Figure 5: Normalized speedup for (a) reconstruction kernel at 75% compression and (b) RoPE-based attention module with an overall 50% compression under different group sizes.
-5 Related Work
-SVD for LLM Compression.
-There are several related works applying SVD for LLM compression.
-An early work (Noach and Goldberg, 2020) simply compresses the weight matrices using standard SVD, resulting in high compression errors.
-To address this, FWSVD (Hsu et al., 2022) incorporates Fisher information to prioritize parameters.
-Additionally, ASVD (Yuan et al., 2023) taken the activation outlier phenomenon in LLM into consideration, and propose and scaling techniques to mitigate the compression error.
-SVD-LLM (Wang et al., 2024) further advances this work by understanding the compression loss from each singular value, improving compression accuracy.
-Unlike previous related SVD works that focus on compressing model weights, Palu focus on using SVD to reduce KV-Cache size.
-KV-Cache Quantization.
-Quantization appears to be an effective compression technique to compress KV-Cache.
-Atom (Zhao et al., 2023) uses a simple per-token quantization scheme to compress KV-Cache.
-WKVQuant (Yue et al., 2024) proposes a two-level quantization scheme to improve quantization accuracy.
-KIVI (Liu et al., 2024) utilizes different quantization schemes, per-channel and per-token, for key and value states respectively.
-KVQuant (Hooper et al., 2024) follow a similar setting but adopts non-uniform quantization and sparse matrix for preserving outliers.
-In Palu, we eliminate outliers with WHT to enhance quantization friendliness and, therefore, only need a simple per-token quantization and achieve a remarkable accuracy.
-6 Conclusion
-We introduce Palu, a novel KV-Cache compression framework that decomposes key and value linear projection weight matrices and caches the compressed latent representations.
-We propose various optimizations, including group-head low-rank decomposition, rank search algorithm, low-rank-aware quantization, matrix fusion, and kernel fusion.
-With these optimizations, Palu can maintain accuracy while achieving significant memory reduction and faster inference.
-Experiments show that Palu can accelerate the end-to-end attention module by up to 1.61$\times$ with 50% KV-Cache compression; when combined with 2-bit quantization, Palu achieves a remarkable 91.25% compression rate, both maintaining strong accuracy.
-Limitations and Future Work
-We introduce a novel KV-Cache compression framework based on low-rank projection.
-However, our experiments cover model variants with up to only 13B parameters due to limited computational resources.
-Future work needs to establish how Palu performs for models with larger sizes (e.g., 70B or beyond).
-Also, we leave for future work the evaluation of end-to-end speedup when combining low-rank compression and quantization.
-It would also be interesting to investigate how to combine Palu with other efficient LLM methods, such as token eviction or weight quantization, to achieve the highest possible acceleration.
-Acknowledgement
-We would like to thank Sandy Kaplan for her valuable feedback and support in refining the writing of this paper.
-Her insights and suggestions significantly contributed to the clarity and quality of the final manuscript.
-Appendix
-Appendix A Quantization Basics
-Quantization techniques use discrete low-bit values to approximate high-precision floating points. The general asymmetric uniform quantization function is defined as:
-$\overline{\mathbf{X}}=\text{clamp}\Big{(}\left\lfloor\frac{\mathbf{X}}{s}\right\rceil+z,0,2^{B}-1),$ (8)
-[…]
-Appendix B Discussion regarding memory usage
-In this work, the experimental results focus on the compression rate of the KV-Cache as a key metric.
-However, it is crucial to consider overall memory savings as a more significant factor.
-For instance, as demonstrated in Sec.
-2.2, a typical compression rate of 30% can lead to an increase in weight size by approximately 40%.
-This increase is calculated under the assumption that $m=n$ and $r=0.7n$, resulting in the equation $\frac{mr+nr}{mn}=1.4$.
-Such an increase indicates substantial extra memory usage.
-This issue primarily arises in J-LRD decomposition schemes, where the projections of all heads are decomposed jointly.
-In contrast, our M-LRD decomposition schemes and optimized G-LRD schemes involve non-square target matrices.
-For example, in the G-LRD scheme with a group size of 4, the target matrix is formed by concatenating the original projection matrices of each attention head in the group.
-In the Llama-2-7b model, with an embedding dimension of 4096 and head dimensions of 128, each projection matrix is 4096x128, resulting in a concatenated matrix of size 4096x512.
-In this case, the dimensions should be considered as $m=8n$.
-Applying the referenced equation $\frac{mr+nr}{mn}$ with $r=0.7n$, we find that $\frac{mr+nr}{mn}=0.7875$, indicating no additional storage cost and, in fact, achieving an additional 21.25% memory savings.
-Furthermore, it is important to highlight that the weights associated with the K and V projections account for only 2 out of 7 linear layers within transformer blocks, comprising merely 16% of the parameters in Llama-2-7b models.
-This limits the overall impact on memory usage.
-Thus, while J-LRD may incur overhead, the M-LRD and G-LRD schemes provide efficient alternatives that do not lead to increased memory usage, making them viable options for practical applications.
-Appendix C Overall Memory Reduction
-In Fig.
-6, we show the total memory usage, including model weights and KV-Cache under different sequence lengths.
-We observe that KV-Cache quickly becomes the memory bottleneck as sequence length grows.
-At a sequence length of 64k, the KV-Cache constitutes 78% of the total memory consumption.
-With 50% low-rank compression, Palu effectively reduces memory usage by 1.7$\times$, and when combined with 2-bit quantization, it further decreases total memory usage by 4.6$\times$.
-Figure 6: Total memory usage for various sequence lengths in Llama2-7B. low-rank compression is at 50% for Palu.
-Appendix D End-to-End Latency Measurements
-We provide the latency evaluation of integrating our optimized attention modules into the Llama-2 models in Table 4.
-Measurement times are expressed in milliseconds (ms).
-Our measurements indicated that Palu achieved significant end-to-end speedup during the decoding phase, especially with longer sequence lengths.
-Specifically, for a sequence length of 32K, the speedups were 1.40x, 1.55x, and 1.71x for gs=4, gs=2, and gs=1, respectively.
-The results demonstrate the effectiveness of our methods and the potential to explore the redundancy in the hidden dimension.
-Table 4: End-to-end decoding latency of Llama-2-7B with Palu measured on A6000.
-Method Group Sequence Length
-Size 4k 8k 16k 32k
-Baseline - 36.07 44.03 69.28 112.75
-J-LRD 32 86.39 103.41 140.85 OOM
-G-LRD 4 32.12 38.89 52.69 80.81
-G-LRD 2 28.82 34.94 47.45 72.89
-M-LRD 1 27.00 32.46 43.56 66.13
-Appendix E Extended Experiments on LongBench and comparisons to quantization baselines
-We conducted extensive experiments using the Mistral-Instruct-7B-v0.2 (Jiang et al., 2023) and LongChat-7b-v1.5-32k (Li et al., 2023) models, both with extended context lengths of 32k tokens.
-We included 16 available English datasets from LongBench to ensure a comprehensive evaluation.
-We report the average score for each type of task separately, as well as the overall average across all 16 tasks.
-This evaluation aimed to broadly assess the effectiveness of Palu in long-context scenarios.
-For reference, we also include comparisons with KIVI.
-Our findings indicate that at a 30% compression level, Palu achieves only a minor average accuracy drop (<1%) compared to the baseline for both models.
-These impressive results are attributed to our effective use of low-rank projection methods to explore redundancy in hidden dimensions.
-Furthermore, by fusing the Hadamard matrix offline to improve quantization friendliness, we were able to quantize the low-rank projected KV-Cache down to 3 bits.
-This quantization process resulted in less than a 1% accuracy degradation, achieved with a straightforward per-token quantization approach.
-When compared to KIVI, our framework demonstrates the ability to harness an additional 30% of redundancy.
-Notably, Palu does not require the complex grouped quantization and mixed-precision techniques employed by KIVI, highlighting the quantization friendliness that Palu offers.
-Table 5: Experiment results on Longbench.
-Average bit widths are estimated for each approach, assuming 10K context length.
-Each column represents the average score for the tasks of each type.
-Multi-QA and Single-QA denote Multi-Document and Single-Document Question Answering, respectively, while Summ refers to Summarization.
-Model Method Avg. Comp. Multi- Single- Summ Few-Shot Code Synthetic Avg.
-Bits Ratio QA QA
+Table 2 showcases the impact of quantization on perplexity and KV-Cache size when combined with Palu.
+With 3-bit quantization, Palu incurs only a slight 0.08 and 0.23 perplexity increase at 30% and 50% low-rank compression rate.
+These demonstrate a minimal accuracy trade-off for significant compression gains compared to the 16-bit baseline.
+Notably, at 2-bit quantization, Palu decisively outperforms the state-of-the-art KVQuant method, reducing perplexity by 1.19 and 0.54, while further slashing memory usage by 30% and 50%.
+These results establish Palu with quantization as a superior KV-Cache compression method.
+4.4 Evaluation on Long Context Datasets
+To access Palu’s ability for long-context scenarios, we evaluate baseline, KIVI and Palu’s accuracy on LongBench (Bai et al., 2023).
+Here we evaluate on Mistral-7B and LongChat-7B models, which have up to 32K context length.
+We report the average score for each task type separately, as well as the overall average across all 16 tasks.
+The results are shown in Table 3.
+As Table 3 indicates, we find that at a 50% low-rank compression level, Palu is relatively difficult to fully preserve accuracy.
+However, at a 30% compression level, Palu achieves only a minor average accuracy drop ($<1\%$) compared to the baseline for both models.
+Furthermore, Palu can quantize the low-rank latent KV-Cache down to 3 bits, with less than 1% further accuracy degradation.
+Overall, Palu maintains a strong 40.77% and 34.33% average accuracy for Mistral-7B and LongChat-7B, with an impressive 7.59x compression ratio.
+When compared to KIVI (Liu et al., 2024b), Palu achieves a similar accuracy, while having an additional 30% compression rate from low-rank.
+Notably, Palu does not require the complex grouped quantization and mixed-precision techniques employed by KIVI, resulting a high inference efficiency (see Sec. 4.5 for details).
+4.5 Latency Evaluation
+In this section, we provide latency and speedup evaluation, using Llama-2-7B as the base model. We measure latency on a single RTX 4090 GPU and compare Palu to the FP16 and KIVI-4-bit baselines.
+We evaluate Palu’s latency at a 50% compression rate, where we set compression rates for key and value to 75% and 25%, respectively.
+This allocation is based on our observations from the rank allocation results (see Appendix G.3 for details).
+For FP16 baseline, we use the default implementation from HuggingFace.
+For KIVI, we use the CUDA kernels from its official repository.
+We do not apply FlashAttention (Dao et al., 2022) for an apple-to-apple comparison to our baselines.
+Due to the small memory capacity of RTX 4090 GPU, we adopt a 4-bit quantization (Frantar et al., 2024) for the weights of all linear layers.
+Our results are the average of 100 runs.
+Table 3: Experiment results on Longbench. Average bit widths are estimated for each approach, assuming 10K context length. Each column represents the average score for the tasks of each type.
+Model Method Avg. Comp. Multi- Single- Summa- Few-Shot Code Synthetic Avg.
+Bits Ratio QA QA rization
 Mistral-7B-v0.2 Baseline 16 1.00x 29.63 36.43 28.10 66.71 54.16 44.87 42.54
@@
 Palu-50% (3 bits) 3 10.6x 18.56 24.14 22.35 58.76 40.50 6.02 29.03
-Table 6: Evaluation integrating LoRA with Palu on Llama2-7B.
+Figure 5: Normalized speedup for both the attention module and end-to-end model decoding.
+Solid lines represent exact measurements, while dashed lines indicate the FP16 baselines are out of memory, and the speedpus are compared the estimated baseline’s latency.
+4.5.1 Speedups of Attention Module and End-to-end Decoding.
+Attention module speedup.
+We compare latency against standard attention without compression or quantization and show the speedups of Palu and KIVI-4-bit in Fig.
+5 (a) and (b) for RoPE-based and non-RoPE attention.
+For RoPE-based attention, we applied our online reconstruction kernel for key and employed offline fusion for value as described in Sec 3.1.2.
+As shown in Fig.
+5 (a), Palu has minimal to no speedup when the sequence length is short, e.g.
+4K.
+However, as sequence length increases, Palu delivers substantial performance gains.
+At 64K input length, Palu achieves a 1.89× speedup over the FP16 baseline when using low-rank projection alone.
+By further applying 4-bit quantization to the Value states, the speedup rises to 2.91$\times$ for the same 64K context length, owing to our optimized low-precision kernel and reduced memory loading times.
+This performance notably surpasses KIVI-4-bit, which only achieves a 1.89× speedup at 64K, hindered by the overheads of its fined-grained group quantization.
+Notably, for RoPE-based attention, Palu-4-bit does not quantize key, as our online reconstruction kernel only supports FP16 precision for now.
+For non-RoPE attention, we apply matrix fusion to both the Key and Value states (Eq.
+6), effectively eliminating all reconstruction overhead.
+At a 64K sequence length with a 50% compression rate, Palu achieves a 2.20× speedup over the FP16 baseline.
+By further applying 4-bit quantization to both the Key and Value states, Palu boosts the speedup to 6.17× for 64K input length.
+These results demonstrate that combining low-rank compression and quantization significantly enhances inference efficiency, particularly in long-context scenarios.
+End-to-end speedup.
+We present the end-to-end speedups in Fig.
+5 (c) and (d), measuring the latency of generating the next token at various input lengths and comparing the results to the FP16 baseline.
+Similar to the attention performance results, Palu shows minimal or no speedup for short sequences but delivers significant acceleration for longer sequences.
+Without quantization, Palu achieves up to 1.71× and 2.05× speedups for RoPE-based and non-RoPE models, respectively.
+With a 50% compression rate, Palu runs up to 32K input length on an RTX 4090 GPU.
+By incorporating 4-bit quantization, Palu handles even longer sequences and delivers 2.59× and 5.53× end-to-end speedups at a 64K sequence length.
+Palu integrated with quantization provides a substantial speed advantage over KIVI-4-bit, which only reaches 1.78× and 1.81× speedups at 32K sequence length for RoPE and non-RoPE scenarios, respectively, and is out-of-memory for longer sequences.
+4.5.2 Kernel for RoPE-Based Attention Score
+In this section, we evaluate the performance of our online reconstruction kernel for RoPE-based attention scores.
+We measure latency from the pre-RoPE query vector to post-GEMV attention score, and compare it with PyTorch’s GEMV, which is used in the baseline attention (see Fig.
+2).
+Figure 6: Speedup of Palu’s attention score kernel with online reconstruction.
+We present speedups for group size 1, 4, and 32 at different sequence lengths in Fig.
+6.
+For gs-32 (J-LRD), the highest accuracy decomposition, the high reconstruction cost causes a significant slowdown across all sequence lengths.
+For gs-1 (M-LRD), our kernel achieves up to a 3.56$\times$ speedup at sequence length 16K, showing strong performance when moderate accuracy loss is acceptable.
+For gs-4 (G-LRD), our kernel reaches up to 1.95× speedup.
+These results emphasize the need to explore various decomposition granularities for better accuracy and speed tradeoffs.
+We also observe that speedup decreases for sequence lengths beyond 16K due to rising reconstruction costs, shifting the online reconstruction from memory- to compute-bound.
+A potential optimization is to quantize the decomposed weight matrices further and leverage high-throughput, low-precision hardware (e.g., INT4 Tensor Cores) for online reconstruction, which we leave for future work.
+Despite the speedup drop at longer lengths, Palu’s overall attention speedup increases with longer input, thanks to matrix fusion on the Value state and the reduced memory footprint.
+5 Related Work
+SVD for LLM Compression.
+Several works have explored using SVD to compress LLMs.
+An early approach (Noach & Goldberg, 2020) applied standard SVD to weight matrices, resulting in significant compression errors.
+FWSVD (Hsu et al., 2022) addressed this by using Fisher information to prioritize parameters, while ASVD (Yuan et al., 2023) considered activation outliers.
+SVD-LLM (Wang et al., 2024) further minimized compression loss for each singular value.
+Unlike these methods, which compress model weights, Palu focuses on reducing KV-Cache size.
+A concurrent work, (Yu et al., 2024), also explores KV-Cache compression using low-rank projection, deriving low-rank matrices from KV-Cache with calibration data and grouping attention heads similarly to Palu’s G-LRD.
+However, it requires LoRA finetuning after decomposition.
+In contrast, Palu directly decomposes the weight matrix, preserving accuracy without finetuning, and introduces additional innovations, including rank search, quantization integration, and optimized GPU kernels.
+KV-Cache Compression.
+Quantization is a widely used technique for compressing KV-Cache.
+Atom (Zhao et al., 2023) applies simple per-token quantization, while WKVQuant (Yue et al., 2024) introduces a two-level scheme to enhance accuracy.
+KIVI (Liu et al., 2024b) uses per-channel and per-token quantization for Keys and Values, combined with fine-grained group quantization at group size 32.
+KVQuant (Hooper et al., 2024) employs a similar setup but incorporates non-uniform quantization and sparse matrices to handle outliers.
+On top of these approaches, GEAR adds a low-rank matrix to compensate for quantization errors.
+In Palu, we leverage low-rank techniques to exploit hidden dimension redundancy and, with simple per-token quantization, achieve outstanding compression results.
+MLA.
+The recently released DeepSeek-V2 model (DeepSeek-AI et al., 2024) introduces the MLA mechanism, which reduces KV-Cache size by down-projecting Key and Value to a low-rank space and reconstructing them to full rank at runtime.
+Although MLA may seem similar to Palu at a high level, particularly with J-LRD, our design and derivation processes are fundamentally different.
+Unlike MLA, a new attention mechanism requiring pre-training, Palu is specifically designed for post-training integration.
+Palu focuses on converting existing models with MHA or GQA to support low-rank compressed KV-Cache, preserving high accuracy while enhancing inference efficiency.
+6 Conclusion
+We introduce Palu, a novel KV-Cache compression framework that decomposes linear projection weight matrices and caches the compressed latent representations.
+We propose various optimizations, including group-head low-rank decomposition, automatic rank allocation algorithm, quantization compatibility enhancement, and customized kernels with operator fusion.
+With these optimizations, Palu can maintain accuracy while achieving significant memory reduction and high inference speedup.
+Experiments show that, with 50% low-rank compression and 4-bit quantization, Palu accelerates RoPE-based attention module by up to 2.91$\times$ and delivers up to 2.2$\times$ end-to-end speedup for the same RoPE-based model, while preserving strong accuracy on various benchmarks.
+Appendix
+Appendix A Quantization Basics
+Quantization techniques use discrete low-bit values to approximate high-precision floating points. The general asymmetric uniform quantization function is defined as:
+$\overline{\mathbf{X}}=\text{clamp}\Big{(}\left\lfloor\frac{\mathbf{X}}{s}\right\rceil+z,0,2^{B}-1),$ (8)
+[…]
+Appendix B Kernel Implementation Details
+Figure 7: Illustration of our fused GPU kernel for computing attention scores with online reconstruction.
+In this figure, $\mathbf{q}$ represents the query vector, $\mathbf{H}$ denotes the low-rank compressed key states, and $\mathbf{B}$ stands for the reconstruction matrices.
+Kernel for attention score calculation with reconstruction.
+The central idea of Palu is to leverage low-rank latent representations to accelerate the attention mechanism by reducing data transfer overhead.
+Instead of working directly with the full-sized key matrix, we store and transfer a compressed low-rank latent representation, denoted as $\mathbf{H}\in\mathbb{R}^{L\times r}$.
+During computation, our custom GPU kernel performs an on-the-fly reconstruction using a reconstruction matrix $\mathbf{B}\in\mathbb{R}^{r\times d_{h}}$, producing a restored key matrix $\mathbf{K}\in\mathbb{R}^{L\times d_{h}}$, where $L$ is the sequence length, $d_{h}$ is the hidden dimension, and $r$ denote the remaining rank after performing low-rank projection.
+The query vector, represented as $\mathbf{q}\in\mathbb{R}^{1\times d_{h}}$, then multiplies with the reconstructed keys to obtain the attention scores.
+To efficiently leverage parallelism, we perform tiling along the sequence length dimension $L$.
+Specifically, we split the sequence into smaller tiles of size $L_{\text{tile}}$, assigning each tile to a dedicated thread block.
+Each thread block independently reconstructs a submatrix $\mathbf{H}_{i}\in\mathbb{R}^{L_{\text{tile}}\times d_{h}}$ from the low-rank latent representation $\mathbf{H}$, then applies the positional embedding using RoPE, and finally performs the matrix-vector multiplication between $\mathbf{q}$ and $\mathbf{H}_{i}$ to produce partial attention scores.
+This design ensures that all intermediate computations, from reconstruction to embedding and final multiplication, remain entirely in on-chip memory (i.e., share memory), thus minimizing high-latency memory access and taking full advantage of the GPU’s parallel processing capabilities to achieve significant speedups.
+Appendix C Discussion regarding memory usage
+In this work, the experimental results focus on the compression rate of the KV-Cache as a key metric.
+However, it is crucial to consider overall memory savings as a more significant factor.
+For instance, as demonstrated in Sec.
+2.2, a typical compression rate of 30% can lead to an increase in weight size by approximately 40%.
+This increase is calculated under the assumption that $m=n$ and $r=0.7n$, resulting in the equation $\frac{mr+nr}{mn}=1.4$.
+Such an increase indicates substantial extra memory usage.
+This issue primarily arises in J-LRD decomposition schemes, where the projections of all heads are decomposed jointly.
+In contrast, our M-LRD decomposition schemes and optimized G-LRD schemes involve non-square target matrices.
+For example, in the G-LRD scheme with a group size of 4, the target matrix is formed by concatenating the original projection matrices of each attention head in the group.
+In the Llama-2-7b model, with an embedding dimension of 4096 and head dimensions of 128, each projection matrix is 4096$\times$128, resulting in a concatenated matrix of size 4096$\times$512.
+In this case, the dimensions should be considered as $m=8n$.
+Applying the referenced equation $\frac{mr+nr}{mn}$ with $r=0.7n$, we find that $\frac{mr+nr}{mn}=0.7875$, indicating no additional storage cost and, in fact, achieving an additional 21.25% memory savings.
+Furthermore, it is important to highlight that the weights associated with the K and V projections account for only 2 out of 7 linear layers within transformer blocks, comprising merely 16% of the parameters in Llama-2-7b models.
+This limits the overall impact on memory usage.
+Thus, while J-LRD may incur overhead, the M-LRD and G-LRD schemes provide efficient alternatives that do not lead to increased memory usage, making them viable options for practical applications.
+Appendix D Overall Memory Reduction
+In Fig.
+8, we show the total memory usage, including model weights and KV-Cache under different sequence lengths.
+We observe that KV-Cache quickly becomes the memory bottleneck as sequence length grows.
+At a sequence length of 64k, the KV-Cache constitutes 78% of the total memory consumption.
+With 50% low-rank compression, Palu effectively reduces memory usage by 1.7$\times$, and when combined with 2-bit quantization, it further decreases total memory usage by 4.6$\times$.
+Figure 8: Total memory usage for various sequence lengths in Llama-2-7B. low-rank compression is at 50% for Palu.
+Table 4: Evaluation integrating LoRA with Palu on Llama-2-7B.
 Comp. Rate Method Zero-Shot Accuracy (%) $\uparrow$
@@
 M-LRD 41.80 70.78 78.02 73.86 43.86 69.22 62.92 -2.12
-Table 7: Zero-shot accuracy of Palu on LLama2 models, with different decomposition strategies at 30% compression rate.
-Model Method Zero-Shot Accuracy (%) $\uparrow$
-OBQA Hella PIQA ARC-e ARC-c Wino Avg. Diff.
-Llama2-7B Baseline 44.20 76.00 78.07 76.30 46.42 69.30 65.05 -
-J-LRD 44.40 76.01 77.75 76.01 45.31 69.38 64.81 -0.24
-G-LRD 45.20 76.07 78.56 75.93 44.97 68.75 64.91 -0.14
-M-LRD 43.20 73.50 77.42 74.16 42.58 68.90 63.29 -1.76
-Llama2-13B Baseline 45.20 79.39 79.11 79.42 49.06 72.38 67.43 -
-J-LRD 45.20 79.58 80.57 79.67 49.40 71.74 67.69 +0.26
-G-LRD 46.60 79.04 78.23 79.42 48.38 71.59 67.21 -0.22
-M-LRD 45.40 78.57 78.78 78.83 47.61 72.06 66.88 -0.55
-Appendix F Integrating Palu with LoRA Finetune
+Appendix E Integrating Palu with LoRA Finetune
 LoRA (Hu et al., 2022) has become one of the most widely used efficient fine-tuning techniques for adapting models to particular tasks or domains with limited data.
@@
 Notably, with LoRA integration, G-LRD shows only a 1.03% accuracy difference compared to J-LRD.
-Appendix G More Results on Zero-shot Accuracy
+Appendix F More Results on Zero-shot Accuracy
 Following up Sec.
-4.1, we further report the zero-shot evaluation results of Palu on the LLama2 family at 30% compression rate.
+4.2, we further report the perplexity and zero-shot evaluation results of Palu on the Llama-2-13B at 50% compression rate.
 As shown in Table.
-7, we observe that Palu achieve nearly no accuracy drops using either J-LRD, G-LRD, or M-LRD.
+5, we observe that Palu achieve competitive accuracy drops around 3% or less across different using either J-LRD, G-LRD, or M-LRD.
 Thus, the users may adopt M-LRD first to optimize the efficiency further.
-Table 8: Ablation study of low-rank decomposition group size on perplexity for the Llama2-7B model at a 50% compression rate using Wikitext-2.
+Table 5: Perplexity and zero-shot accuracy of Palu, with different decomposition strategies at 50%
+Model Method Perplexity $\downarrow$ Zero-Shot Accuracy (%) $\uparrow$
+Wiki2 C4 OBQA Hella PIQA ARC-e ARC-c Wino Avg.
+Llama-2-13B Baseline 4.88 6.70 45.20 79.39 79.11 79.42 49.06 72.38 67.43
+J-LRD 4.97 6.92 46.40 79.48 78.62 79.29 49.91 70.56 67.38
+G-LRD 5.31 7.76 45.60 77.29 77.42 76.05 45.99 72.45 65.80
+M-LRD 5.65 8.34 43.20 74.34 77.53 75.76 45.39 68.98 64.20
+Table 6: Ablation study of low-rank decomposition group size on perplexity for the Llama-2-7B model at a 50% compression rate using Wikitext-2.
 Method Group Size Perplexity
@@
 M-LRD 1 6.81
-Appendix H Ablation Study
-H.1 Influence of Different Group Size
+Appendix G Ablation Study
+G.1 Influence of Different Group Size
 Since our proposed G-LRD method allows for balancing performance and efficiency by adjusting the group size, we conducted an ablation study on group size.
 As seen in Table 6, as the group size increases, the amount of shared information also increases, leading to improved performance.
-H.2 Influence of Walsh-Hadamard Transform
+G.2 Influence of Walsh-Hadamard Transform
 We conduct the ablation study to profile the benefits of applying the Walsh-Hadamard Transform (WHT).
@@
 3.4 for more details.
-H.3 Automatic Rank Allocation vs. Uniform Rank Allocation
+G.3 Automatic Rank Allocation vs. Uniform Rank Allocation
 Table 8 presents the ablation study on the impact of different rank allocation schemes on the model’s accuracy.
@@
 The results clearly demonstrate a non-uniform allocation result.
-Specifically, we observe that the value is generally allocated a higher rank compared to the key.
+Specifically, we observe that the value is generally allocated a higher rank than the key.
 Additionally, the first half of the layers are assigned higher ranks, indicating their greater importance in preserving model performance.
@@
 Same as Sec.
-4.2, we use the WikiText-2 with sequence length set to 4096 as the evaluation benchmark.
+4.3, we use the WikiText-2 with sequence length set to 4096 as the evaluation benchmark.
 Method Wikitext-2 PPL $\downarrow$
@@
       + 2-bits w Hadamard 6.41 (4.17$\downarrow$)
-Figure 7: Visualization of layer-wise low-rank compression rate on Llama2-7B with 50% of overall compression rate.
+Figure 9: Visualization of layer-wise low-rank compression rate on Llama-2-7B with 50% of overall compression rate.
 Here, compression rates (i.e., rank) are allocated using the proposed Fisher Information-based automated rank allocation algorithm.
-Table 10: Ablation study on w/ and w/o rank search. We use Llama2-7b and Wikitext-2 with sequence length 2048 as the benchmark.
+Table 8: Ablation study on w/ and w/o rank search. We use Llama-2-7b and Wikitext-2 with sequence length 2048 as the benchmark.
 Rate=30% Rate=50% Rate=70%
@@
 Automatic (ours) 5.62 (0.72$\downarrow$) 6.02 (1.36$\downarrow$) 8.59 (2.18$\downarrow$)
-Appendix I Experiment Details
-I.1 Zero-Shot Evaluation Details
+Appendix H Experiment Details
+H.1 Zero-Shot Evaluation Details
 We selected six zero-shot tasks from the LM-eval benchmark to evaluate Palu:
 We report accuracy for WinoGrande, PIQA, and ARC-Easy, and accuracy normalized by sequence length (acc_norm) for HellaSwag and ARC-Challenge.
-I.2 LongBench Evaluation Details
+H.2 LongBench Evaluation Details
 For the LongBench evaluation in the manuscript, we selected eight tasks from four subgroups, ensuring a comprehensive evaluation of Palu. The tasks and their corresponding metrics are detailed below:
-During evaluation, we set the maximum sequence length to 8192 for the Mistral model and 15500 for Vicuna.
+During the evaluation, we set the maximum sequence length to 31500 for both the Mistral and LongChat model.